### v1

# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Simulation, and Real-Vehicle Experiment

Abstract
With the broader usage and highly successful development of Large Language Models (LLMs), there has been a growth of interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning ability, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to language interaction and decision-making. In this paper, we first introduce novel concepts and approaches to designing LLMs for autonomous driving (LLM4AD). Then, we propose a comprehensive benchmark for evaluating the instruction-following abilities of LLMs within the autonomous driving domain. Furthermore, we conduct a series of experiments on both simulation and real-world vehicle platforms, thoroughly evaluating the performance and potential of our LLM4AD systems. Our research highlights the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making.
Index Terms:
Large Language Models; Autonomous Driving; Human-Centric AI; Human-Autonomy Teaming; Embodied AI.
I Introduction
I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from the classic nature language process (NLP) tasks like document modification and information extraction to the new emerging scenarios such as LLM-based agents and LLMs for evaluation [zhao2023survey]. One particular application that our study focuses on, among many other domains, is adopting LLMs for autonomous driving (LLM4AD). In this area, various advanced emerging LLM-based algorithms and technologies are continually enhancing the capabilities of autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency forward. LLMs can contribute to autonomous systems from high-level decision-making processes to meticulous low-level control. On the high-level side, LLMs can actively engage in adjusting driving mode or decision-making process [cui_drive_2024]. Imagine the scenario in which you are seated in an autonomous vehicle, where all you need to do is abstractly express your feelings like “I do not want my friends to wait for me.” The LLM4AD system then interprets your emotions and adjusts the vehicle’s control strategies accordingly, aligning with your current driving mood or preferences. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret humans’ intentions only from some vague expressions [cui2023large]. Besides, LLMs-based systems have significant potential for achieving personalization in autonomous driving due to their continuous learning capacity. This capability enables it to continuously adapt to individuals’ preferences and desires, improving the driving experience according to different users. In addition, LLMs also hold the potential to develop a knowledge-driven system capable of driving like an expert, which can accumulate experience through continuous driving [fu_drive_2024].
On the other hand, on the low-level side, LLMs can also play a crucial role in the tuning and control process. LLMs have demonstrated the ability to analyze specific scenarios and convert gathered information into mathematical representations to guide low-level controllers [sha_languagempc_2023]. Additionally, LLMs can receive input data from the controller and provide performance updates to aid humans in analyzing the control loop’s effectiveness [s23135899]. They can potentially suggest improvements or detect issues to enhance overall performance.
The emerging developments of LLM4AD systems raise a crucial question for all related researchers: Why have LLMs become incredibly popular in autonomous driving? What are the advantages of LLM4AD systems over those without LLM integration? Based on current research and developments, we summarize several compelling advantages of integrating LLMs as follows:
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Humans can express abstract commands and feelings, while the LLMs can accurately capture the intentions behind human expressions.
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources such as traffic laws and accident reports, allowing generated decisions to guarantee safety and follow local regulations.
  • Zero-Shot and Few-Shot Planning: The zero-shot generalization ability enables LLMs to perform tasks they have not been trained on before [kojima_large_2022]. This means LLM-based autonomous systems can handle uncommon situations with minimal or no prior experience, enabling them to confidently navigate through ‘corner cases.’
  • Continuous Learning and Personalization: LLMs learn and adapt continuously, providing the ability to follow individual user preferences and improve the driving experience over time.
  • Interpretability and Trust: LLMs can explain their decisions in natural language, improving trust and understanding between autonomous systems and their users.
Despite the numerous advantages above, LLMs still have some limitations and potential issues. The most obvious one is sometimes they are inapplicable to real-time tasks in autonomous driving. Specifically, LLMs typically require seconds to process textual information. This delay can pose significant safety concerns and risks in situations requiring immediate, real-time decision-making. Additionally, addressing hallucinations is also a main challenge in LLM4AD systems. “Hallucination” refers to instances where LLMs generate output that is factually incorrect, nonsensical, or unrelated to the input prompt. As autonomous driving is a safety-critical task, hallucinations can introduce significant safety and reliability concerns and potentially undermine trust in LLM4AD systems. Furthermore, data privacy and security are other potential issues since LLMs collect and process vast amounts of textual data, including potentially sensitive information about their surroundings, passengers, and driving preferences.
I-B Overview of This Article
Our paper presents a comprehensive study on integrating LLMs into autonomous driving systems, from the proposed concept and the creation of an open dataset to the implementation in simulations and real-vehicle experiments. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III develops an open benchmark designed to evaluate the performance of LLMs in various autonomous driving tasks, including dataset, evaluation metrics, and baseline model. Sec. IV validates the proposed concepts through extensive simulations, exploring a key perspective: learning from human feedback. LABEL:sec:real-vehicle presents the implementation of our framework on a real-vehicle platform and discusses the results of a series of experiments conducted to validate the system’s driving and personalization performance. LABEL:sec:discussion analyzes the current limitations and outlines promising future research directions for leveraging LLMs in autonomous driving applications. Finally, LABEL:sec:conclusion summarizes the key contributions and findings of this work.
Figure 1: Our proposed framework is an LLM-driven autonomous driving system. Within our framework, LLMs receive inputs including system messages, human instructions, situation descriptions, and the historical memory of the specific user. The outputs consist of two main components: a generated program and reasoned thoughts. The generated program is subsequently executed within the executor, interacting with the environment.
TABLE I: Symbols and their meanings used in this document.
Symbol Meaning
LLM Large Language Model
$I$ Instruction
$C$ Contextual (Environment) Information
$S$ System Message
$H$ History Data
$F$ Feedback
$R$ Reasoning Thoughts
$P$ Generated Program
$P^{\prime}$ Improved Generated Program
LMP Language Model Program
$\ell$ LLM backbone
API Application Programming Interface
TTC Time to Collision
SV Speed Variation
TE Time Efficiency
II Concept of LLM4AD
In this section, we introduce a perspective where LLMs play the role of the decision-making “brain” within the autonomous driving system. Within our framework, LLMs do not directly affect perception or localization modules, which serve as the vehicle’s “eyes;” instead, they utilize outputs from these modules as references to guide the high-level decision-making process. By receiving processed data from them, LLMs can enhance informed decision-making, leading to significant improvements in the performance of autonomous vehicles. At the downstream, the vehicle’s control module serves as its “hands,” executing the driving policy obtained from the LLM-based decision-making process. The symbols we used in the rest of the paper are listed in Tab. I.
The overall LLM4AD framework is shown in Fig. 1. In our framework, humans provide instructions and evaluations, where the instructions $I$ and evaluations $F$ along with historical memories $H$, system messages $S$, and context information $C$ serve as inputs to LLMs. The memory module stores corresponding historical interactions between humans and vehicles $H$ for different users. Upon receiving these inputs, LLMs engage in reasoning and produce outputs, including generated Language Model Programs (LMPs) $P$ and reasoning thoughts $R$. The generated LMP is sent to the executor for execution in the environment, while the reasoning thoughts help LLMs generate more reasonable driving policies. It is important to note that this is a general concept, where specific implementations may vary across different applications. However, all applications follow the principles and framework of this concept.
II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are directly inputted to LLMs by humans in natural language. $I$ includes the humans’ desired needs for the autonomous agent, while human evaluations $F$ are the feedback on the effectiveness of these driving policies.
II-B System Message
The system message $S$ is a way to provide instructions or context to the LLM4AD system at the start of a conversation or task. System messages $S$ allow developers to clearly define the role, constraints, and objectives that LLM4AD systems should follow when engaging with users or working on particular types of tasks.
System messages $S$ in autonomous driving tasks act like a set of high-level guidelines or rules that frame how LLMs should operate and make decisions. These high-level system messages are crucial in defining the basic driving logic, including tasks such as task definition, adherence to traffic rules, description of decision states, and overall goals or metrics to optimize for. They serve as a foundational framework guiding the behavior and decision-making processes of autonomous vehicles on the road. Without a well-crafted system message, LLMs could make incorrect assumptions or unintended strategies.
II-C Situation Descriptor
A situation descriptor transfers the current driving context $C$ into textual descriptions. The situation descriptor follows a predefined structure and employs natural language to provide a detailed textual description of the current driving scenario. It aims to provide LLMs with situational awareness and a comprehensive representation of the current driving scenario, allowing them to make appropriate decisions in the current traffic situation. This could include descriptive statements like “You are positioned in the leftmost lane of a two-way highway” or “A vehicle is located 50 meters ahead of your current position.” The descriptor translate s complex spatial and temporal relationships between different road users and ego vehicles intuitively into the natural language format, allowing context information to be reasoned by LLMs.
II-D History Memory and Memory Module
The memory module stores profiles from various users to enhance driving experiences for all users. Whenever a human user utilizes the LLM4AD system, the system logs the relevant historical interactions $H$ associated with that user. Subsequently, the historical data for the current human user is transmitted to LLMs as input. This particular history interaction $H$ serves as a reference point for the preferences of the current users, therefore guiding the system to improve user experience. After each trip, the interaction data will be updated in the corresponding profile in the memory module. The specifics of the interaction data vary for each system and will be elaborated in their respective sections.
II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history memories (interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human evaluation $F$) and generate the textual outputs (LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations. Chain-of-thought presents a structured sequence of reasoning examples to LLMs, bridging existing knowledge gaps. By offering a series of logical and connected steps, LLMs can perform well in complex driving scenarios more effectively.
II-F Generated Program
Inspired by the concept of “Code as Policy [liang_code_2023],” one of the key outputs from LLMs is the generated LMPs $P$ consisting of executable codes. These codes are used to affect the driving behavior of the ego agent within the environment, and they not only have the ability to generalize to new natural language commands but can also provide precise numerical values such as velocities based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
II-G Output Thought
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is ‘hurry,’ I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated programs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.
II-H Executor
The executor acts as the bridge between LLMs’ textual outputs and the current autonomous driving policy. It takes the generated LMPs $P$ from LLMs and executes it in the corresponding environment. This allows the codes to interact with the current status of the ego vehicle and enables the generated programs to deploy their intended driving behaviors in the real or simulated environment. In our scope, different autonomous driving systems will have some differences in how those codes are executed by their respective executors.
III An Benchmark Dataset of LLM4AD
In this section, we introduce LaMPilot-Bench, the first benchmark of its kind for evaluating the instruction-following capabilities of LLM-based agents in autonomous driving. LaMPilot-Bench consists of three key components: a simulator, a dataset, and an evaluator.
III-A Simulator
The LaMPilot-Bench simulator is built upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making. HighwayEnv offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
III-B Dataset
The LaMPilot dataset consists of 4,900 semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes:
  • An instruction $I$: a high-level task description.
  • An initial state: used to initialize the simulator.
  • Goal state criteria: aligned with the instruction $I$.
The dataset covers diverse driving scenarios, as shown in Tab. II. For each driving scenario, LaMPilot includes various situations. Taking turning scenarios as an example, the diversity is reflected in several variables, such as the ego vehicle’s initial position and state, the specific task (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed. The dataset also contains a variety of instructions reflecting realistic in-cabin human commands categorized by maneuver types (e.g., routing, lane changing, overtaking) and scenario types (highway and intersection). Fig. 2 shows the distribution of the first four words in the instructions, while main statistics are provided in Tab. II.
TABLE II: Main statistics of the LaMPilot-Bench dataset.
Statistic Value
Total scenarios 4,900 (100%)
Distance-related 1,200 (24.5%)
Speed-related 1,200 (24.5%)
Pulling over 200 (4.1%)
Routing 1,500 (30.2%)
Lane changing 400 (8.2%)
Overtaking 400 (8.2%)
Highway 3,400 (69.4%)
Intersection 1,500 (30.6%)
Average instruction length 7.6 words
Maximum instruction length 14 words
Minimum instruction length 2 words
Figure 2: Distribution of instructions in the LaMPilot-Bench dataset.
III-C Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) is used to measure the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 00), the TTC ($\tau_{i}$) with vehicle $i$ is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
$\tau_{i}=-\frac{(\mathbf{p}_{0}-\mathbf{p}_{i})\cdot(\mathbf{v}_{0}-\mathbf{v}_{i})}{\|\mathbf{v}_{0}-\mathbf{v}_{i}\|^{2}}$ (1)
The minimum positive TTC value $\tau_{\min}$ among all $n$ vehicles over all time steps $t$ (up to task completion time $T$) represents the nearest potential collision:
$\tau_{\min}=\min_{1\leq i\leq n,,1\leq t\leq T}\tau^{t}_{i}$ (2)
TTC scores are empirically assigned based on a 2-second safety margin, with values above 2 seconds considered safe and given a score of 100:
$\operatorname{TTC}=\begin{cases}100&\text{if $\tau_{\min}>2$}\\ \max(0,100-\frac{1}{\tau_{\min}})&\text{if $0<\tau_{\min}\leq 2$}\\ 0&\text{if $\tau_{\min}\leq 0$}\end{cases}$ (3)
Speed variance (SV) is considered as another safety metric [garber_factors_1989]. The ego vehicle’s speed standard deviation $\sigma_{0}$ is calculated as:
$\sigma_{0}=\sqrt{\frac{\sum_{t=1}^{T}{(\|\mathbf{v}^{t}_{0}\|-\mu_{0})^{2}}}{T}},$ (4)
where $\mu_{0}$ is the mean speed:
$\mu_{0}=\frac{\sum_{t=1}^{T}{\|\mathbf{v}^{t}_{0}\|}}{T}.$ (5)
The SV score is defined as:
$\operatorname{SV}=\max(0,100\cdot(1-{\sigma_{0}}/{\sigma_{\operatorname{safe}}})),$ (6)
where $\sigma_{\operatorname{safe}}$ is the maximum safe speed deviation, determined empirically. The time efficiency (TE) score evaluates the policy’s ability to complete the task within a predefined time limit $T_{\operatorname{limit}}$:
$\operatorname{TE}=\max(0,100\cdot(1-{T}/{T_{\operatorname{limit}}}))$ (7)
A TE score closer to 100 indicates a more efficient policy.
A task is considered successfully completed when the agent achieves the objectives specified in the instruction while maintaining safety (i.e., avoiding collisions) and efficiency (i.e., finishing within the allotted time). For example, a lane change task is completed when the vehicle is in the target lane with its heading aligned with the lane’s direction within a specified threshold. The final score aggregates all individual metrics, weighted according to their importance:
$\operatorname{Score}=W_{\operatorname{TTC}}\cdot\operatorname{TTC}+W_{\operatorname{SV}}\cdot\operatorname{SV}+W_{\operatorname{TE}}\cdot\operatorname{TE}$ (8)
TABLE III: Comparison of baselines on LaMPilot-Bench under different learning settings.
Method Learning Collision ($\downarrow$) Completion ()($\uparrow$)( ↑ ) TTC Score ()($\uparrow$)( ↑ ) SV Score ()($\uparrow$)( ↑ ) TE Score ()($\uparrow$)( ↑ ) Driving Score ()($\uparrow$)( ↑ )
IDM N/A 0.0% 20.4% 92.8 80.9 71.0 17.3
MOBIL 0.0% 15.3% 82.8 85.9 46.7 11.7
Human Driver 0.0% 98.0% 93.4 86.3 81.3 84.6
Llama 2 0-Shot 0.0% 20.4% 92.8 81.4 68.9 17.3
PaLM 2 3.1% 35.7% 78.6 76.5 78.8 12.8
ChatGPT 1.0% 40.8% 83.9 75.5 74.0 27.8
GPT-4 4.1% 72.4% 61.3 70.5 74.0 28.5
GPT-4 Turbo 3.1% 74.5% 73.2 70.1 73.9 39.1
Llama 2 3-Shot 1.0% (+1.0) 63.3% (+42.9) 68.3 (-24.5) 71.4 (-10.0) 73.8 (+4.9) 39.6 (+22.3)
PaLM 2 3.1% (0.0) 71.4% (+35.7) 73.4 (-5.2) 69.7 (-6.8) 72.0 (-6.8) 36.6 (+23.8)
ChatGPT 4.1% (+3.1) 83.7% (+42.9) 70.9 (-13.0) 73.7 (-1.8) 77.7 (+3.7) 41.2 (+13.4)
GPT-4 1.0% (-3.1) 84.7% (+12.3) 69.4 (+8.1) 72.0 (+1.5) 76.7 (+2.7) 55.9 (+27.4)
GPT-4 Turbo 1.0% (-2.1) 80.6% (+6.1) 69.1 (-4.1) 71.5 (+1.4) 74.9 (+1.0) 52.6 (+13.5)
Llama 2 Human Feedback 0.9% (-0.1) 32.7% (-30.6) 83.9 (+15.6) 74.6 (+3.2) 74.6 (+0.8) 21.4 (-18.2)
PaLM 2 0.0% (-3.1) 45.5% (-25.9) 81.1 (+7.7) 74.4 (+4.7) 77.6 (+5.6) 35.6 (-1.0)
ChatGPT 0.9% (-3.2) 80.9% (-2.8) 70.2 (-0.7) 71.7 (-2.0) 77.5 (-0.2) 54.2 (+13.0)
GPT-4 0.9% (-0.1) 92.7% (+8.0) 67.6 (-1.8) 72.2 (+0.2) 76.8 (+0.1) 64.0 (+8.1)
GPT-4 Turbo 0.9% (-0.1) 87.3% (+6.7) 74.1 (+5.0) 73.4 (+1.9) 75.5 (+0.6) 60.5 (+7.9)
III-D Experiment Setup
The experiments are designed to establish baselines for the LaMPilot-Bench. The main objectives are to assess the performance of LLM-based agents in interpreting human instructions within driving contexts and to evaluate the capability of LLMs to generate code for motion planning using provided functional primitives.
We conduct benchmarking using various state-of-the-art large language models, including both open-source and proprietary solutions [^1] OpenAI API accessed in October and November 2023.:
  • Llama 2 [touvron_llama_2023] (llama-2-70b-chat)
  • PaLM 2 [anil_palm_2023] (code-bison)
  • ChatGPT [openai_introducing_2023] (gpt-3.5-turbo)
  • GPT-4 [openai_gpt-4_2023] (gpt-4)
  • GPT-4-Turbo [openai_gpt-4-turbo_2023] (gpt-4-1106-preview)
We set a 60-second time limit for each scenario. If a task is not completed within this time frame, the test case is considered a failure, and the simulation is terminated, resulting in a driving score of 0. For successful cases, the driving score is calculated as follows:
$\operatorname{Driving\ Score}=\frac{\alpha}{N_{s}}\sum_{i=1}^{N_{s}}\operatorname{Score}_{i}-\beta\cdot P_{\operatorname{collision}}$ (9)
Here, $N_{s}$ represents the number of successful test cases, $\alpha$ is the success rate (ranging from 0 to 1), $\beta$ is the collision rate (also between 0 and 1), and $P_{\text{collision}}$ is the penalty factor for collisions (set at 500 in our experiments). The $\text{Score}_{i}$ for each individual sample is calculated based on Eq. 8, with weights $W_{\text{TTC}}$, $W_{\text{SV}}$, and $W_{\text{TE}}$ set to 0.5, 0.3, and 0.2, respectively. The driving score provides a comprehensive assessment of the agent’s performance in LaMPilot-Bench.
III-E Baselines
III-E1 Heuristic Baselines
In our experiments, we employ two rule-based baseline policies: the Intelligent Driver Model (IDM)[treiber_congested_2000] and the Minimizing Overall Braking Induced by Lane Changes (MOBIL) principle[kesting_general_2007]. IDM describes a rule for updating the acceleration of a vehicle to avoid collisions based on the proximity and relative velocity of the vehicle to the object directly in front. MOBIL is an extension of IDM where a lane change is executed if the prospective new lane offers a more favorable driving scenario and the maneuver can be conducted safely. These baselines can be considered as operating on random chance, as the policy is independent of user instructions but follows predefined rules. Additionally, we include a human performance baseline, where a licensed human driver controls the vehicle in the simulation using arrow keys on the keyboard, following the commands and visuals displayed. This baseline provides a reference for human-level performance on LaMPilot-Bench.
III-E2 Zero-Shot and Few-Shot Baselines
We also include the zero-shot and few-shot CoT prompting [wei_chain--thought_2022] as baseline methods. The zero-shot CoT baseline leverages the prompt “How to complete the task step by step.” In the few-shot setting, we follow the standard practice [brown_language_2020] by including 3 human-written code exemplars before the test instance. These in-context examples help the model adapt to the tasks in LaMPilot-Bench. The in-context examples are created by a programmer proficient in Python. They are provided with API descriptions and are allowed to write and test their code using the non-test scenarios. The same set of three examples are used for all test cases. The API Docs are also the same across all test cases for both zero-shot and few-shot settings.
III-E3 Human Feedback Baselines
LLMs have demonstrated proficiency in generating coherent solutions across various tasks without requiring additional fine-tuning. However, when it comes to code generation, especially for complex scenarios, they can produce suboptimal results. The autoregressive nature of these models poses a significant challenge, as tokens generated early in a sequence cannot be modified in the same iteration. This constraint limits the models’ ability to refine initial responses, potentially impacting the effectiveness of the generated code [wang_voyager_2023].
To address these challenges and enhance the performance of LLMs on tasks in LaMPilot-Bench, we include the human-in-the-loop baseline introduced in LABEL:sec:human-in-the-loop-learning, where the LLMs serve not only as planners but also as engines for incorporating human feedback. After an LLM generates a policy program ($P$), it can integrate human feedback ($F$) with context-specific guidance, enabling the LLM to refine its output.
This approach does not involve fine-tuning or training a new LLM. Instead, we leverage the RAG approach [lewis_retrieval-augmented_2020] and introduce a code repository module, which is implemented with the Chroma [chroma_ai-native_2023] vector database. This repository allows for the storage and retrieval of effective code snippets to be used in similar situations. Following Voyager [wang_voyager_2023], we use LLM-generated function descriptions, which are then converted into vectors to serve as database keys. These keys are paired with the corresponding function codes as their values. If the human is satisfied with the policy, the code is committed to the repository for future reference. The human feedback learning process is conducted on the non-test set, involving interaction with 100 random samples for each method to ensure fair comparisons.
III-F Quantitative Results
In this section, we present the experimental results on the LaMPilot-Bench, summarizing the performance of various methods, including heuristic baselines, zero-shot and three-shot baselines, and human feedback baselines. These results are shown in Tab. III.
Rule-Based Methods
We first evaluate the performance of rule-based approaches, specifically the IDM and MOBIL algorithms. Both methods achieve a zero collision rate in LaMPilot-Bench. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
LLM-Based Agents
We assess the effectiveness of off-the-shelf LLMs to reason in the context of autonomous driving and following human instructions. In the zero-shot setting, GPT models and PaLM 2, given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion. This improvement can be attributed to the LLMs’ ability to understand and generate code based on natural language instructions, leveraging their extensive pretraining on diverse data sources. However, the zero-shot setting also results in an increase in the collision rate (1%-4%), indicating that the generated code policies do not fully capture the complexities of driving tasks.
When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting. This improvement demonstrates the effectiveness of few-shot learning in adapting LLMs to specific task domains, as the exemplar code provides a clearer understanding of the desired behavior and API usage patterns.
The integration of GPT models with human-in-the-loop learning further enhances their performance in driving tasks. Notably, GPT-4 achieves the highest driving score of $64.0$. This improvement highlights the great potential of LLMs in following instructions in the driving context when combined with human feedback. The human feedback loop allows for iterative refinement of the generated policies, enabling the LLMs to learn from their mistakes and continuously incorporate human knowledge to generate more effective driving strategies.
III-G Conclusion
In this section, we introduced LaMPilot-Bench, a benchmark for evaluating the instruction-following capabilities of LLM-based agents in AD. Our experiments demonstrated that off-the-shelf LLMs can generate code policies for driving tasks based on human instructions. However, the notable collision rate indicates the need for further research to fully capture the complexities and safety requirements of real-world driving scenarios. As LLMs continue to evolve and integrate with more modalities, we envision a future where they can significantly contribute to the development of safe, efficient, and user-friendly autonomous vehicles.
IV Simulation Studies of LLM4AD
TABLE IV: An example that illustrates the human-feedback learning process.