### v3

# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

Abstract
With the broader usage and highly successful development of Large Language Models (LLMs), there has been a growth of interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning ability, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to language interaction and decision-making. In this paper, we first introduce the novel concept of designing LLMs for autonomous driving (LLM4AD). Then, we propose a comprehensive benchmark for evaluating the instruction-following abilities of LLM4AD in simulation. Furthermore, we conduct a series of experiments on real-world vehicle platforms, thoroughly evaluating the performance and potential of our LLM4AD systems. Finally, we envision the main challenges of LLM4AD, including latency, deployment, security and privacy, safety, trust and transparency, and personalization. Our research highlights the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making.
Index Terms:
Large Language Models; Autonomous Driving; Human-Centric AI; Human-Autonomy Teaming; Embodied AI.
I Introduction
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=11em,font=,, where level=3text width=18em,font=,, [ Large Language Models for Autonomous Driving, ver [ Concept of LLM4AD ] [ Benchmark and Simulation [ LaMPilot-Bench, leaf ] [ CARLA Leaderboard 1.0 Benchmark, leaf ] [ Baselines and Results, leaf ] [ Summary, leaf ] ] [ Real-Vehicle Experiments:
On-Cloud LLM for Personalized
Decision-Making [ Talk2Drive Framework, leaf ] [ Autonomous Driving System, leaf ] [ Experiment Setup, leaf ] [ Experiment Results, leaf ] ] [ Real-Vehicle Experiments:
On-Edge VLM for Personalized
Motion Control [ Problem Statement, leaf ] [ Our Framework, leaf ] [ Real-World Experiment, leaf ] ] [ Challenges [ Latency, leaf ] [ Security and Privacy, leaf ] [ Development, leaf ] [ Safety, leaf ] [ Personalization, leaf ] [ Trust and Transparency, leaf ] ] ]
Figure 1: A condensed overview of this paper.
I-A Background
The recent development of Large Language Models (LLMs) has enhanced numerous cutting-edge technologies, introducing significant advancements across a broad spectrum of applications [zhao2023survey, wang2023survey]. Their applications span from the classic nature language process (NLP) tasks like document modification and information extraction to the new emerging scenarios such as LLM-based agents and LLMs for evaluation [zhao2023survey]. One particular application that our study focuses on, among many other domains, is adopting LLMs for autonomous driving (LLM4AD). In this area, various advanced emerging LLM-based algorithms and technologies are continually enhancing the capabilities of autonomous driving technology, leveraging the potential of LLMs to drive innovation and efficiency forward. LLMs can contribute to autonomous systems from high-level decision-making processes to meticulous low-level control. On the high-level side, LLMs can actively engage in adjusting driving mode or decision-making process [cui_drive_2024]. Imagine the scenario in which you are seated in an autonomous vehicle, where all you need to do is abstractly express your feelings like “I do not want my friends to wait for me.” The LLM4AD system then interprets your emotions and adjusts the vehicle’s control strategies accordingly, aligning with your current driving mood or preferences. By contrast, non-LLM-based systems lack the ability to accurately comprehend or interpret humans’ intentions only from some vague expressions [cui2023large]. Besides, LLMs-based systems have significant potential for achieving personalization in autonomous driving due to their continuous learning capacity. This capability enables it to continuously adapt to individuals’ preferences and desires, improving the driving experience according to different users. In addition, LLMs also hold the potential to develop a knowledge-driven system capable of driving like an expert, which can accumulate experience through continuous driving [fu_drive_2024].
On the other hand, on the low-level side, LLMs can also play a crucial role in the tuning and control process. LLMs have demonstrated the ability to analyze specific scenarios and convert gathered information into mathematical representations to guide low-level controllers [sha_languagempc_2023]. Additionally, LLMs can receive input data from the controller and provide performance updates to aid humans in analyzing the control loop’s effectiveness [s23135899]. They can potentially suggest improvements or detect issues to enhance overall performance.
The emerging developments of LLM4AD systems raise a crucial question for all related researchers: Why have LLMs become incredibly popular in autonomous driving? What are the advantages of LLM4AD systems over those without LLM integration? Based on current research and developments, we summarize several compelling advantages of integrating LLMs as follows:
  • Intuitive Language Interaction: LLMs enable intuitive communication between humans and vehicles. Humans can express abstract commands and feelings, while the LLMs can accurately capture the intentions behind human expressions.
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources such as traffic rules and accident reports, allowing generated decisions to guarantee safety and follow local regulations.
  • Zero-Shot and Few-Shot Planning: The zero-shot generalization ability enables LLMs to perform tasks they have not been trained on before [kojima_large_2022]. This means LLM-based autonomous systems can handle uncommon situations with minimal or no prior experience, enabling them to confidently navigate through “corner cases.”
  • Continuous Learning and Personalization: LLMs learn and adapt continuously, providing the ability to follow individual user preferences and improve the driving experience over time.
  • Interpretability and Trust: LLMs can explain their decisions in natural language, improving trust and understanding between autonomous systems and their users.
Despite the numerous advantages above, LLMs still have some limitations and potential issues. The most obvious one is sometimes they are inapplicable to real-time tasks in autonomous driving. Specifically, LLMs typically require seconds to process textual information. This delay can pose significant safety concerns and risks in situations requiring immediate, real-time decision-making. Additionally, how to avoid the negative effect provided by hallucinations is also a main challenge in LLM4AD systems. “Hallucination” refers to instances where LLMs generate output that is factually incorrect, nonsensical, or unrelated to the input prompt. As autonomous driving is a safety-critical task, hallucinations can introduce significant safety and reliability concerns and potentially undermine trust in LLM4AD systems. Furthermore, data privacy and security are other potential issues since LLMs collect and process vast amounts of textual data, including potentially sensitive information about their surroundings, passengers, and driving preferences.
I-B Overview of This Article
Fig. 1 provides an overview of the reminder structure of this paper. Our paper presents a comprehensive study on integrating LLMs into autonomous driving systems, from the proposed concept and the creation of an open dataset to the implementation in simulations and real-vehicle experiments. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III develops an open benchmark designed to evaluate the performance of LLMs in various autonomous driving tasks, which includes a dataset, evaluation metrics, and baseline models. Sec. IV and Sec. V present the implementation of our framework on a real-vehicle platform and discuss the results of a series of experiments conducted to validate the system’s driving and personalization performance. Sec. VI analyzes the current challenges and outlines promising future research directions for leveraging LLMs in autonomous driving applications. Finally, Sec. VII summarizes the key contributions and findings of this work.
Figure 2: The conceptual illustration of the proposed LLM4AD system.
TABLE I: Symbols and their meanings used in this document.
Symbol Meaning
LLM Large Language Model
$I$ Instruction
$C$ Contextual (Environment) Information
$S$ System Message
$H$ History Data
$F$ Feedback
$R$ Reasoning Thoughts
$P$ Generated Program/Policy
$V$ Visual Input
$D$ Scene Description
$A$ Executed Action
LMP Language Model Program
$\ell$ LLM backbone
API Application Programming Interface
TTC Time to Collision
SV Speed Variation
TE Time Efficiency
II Concept of LLM4AD
In this section, we introduce a perspective where LLMs play the role of the decision-making “brain” within the autonomous driving system. Within our framework, LLMs do not directly affect perception or localization modules, which serve as the vehicle’s “eyes;” instead, they utilize outputs from these modules as references to guide the high-level decision-making process. By receiving processed data from them, LLMs can enhance informed decision-making, leading to significant improvements in the performance of autonomous vehicles. At the downstream, the vehicle’s control module serves as its “hands,” executing the driving policy obtained from the LLM-based decision-making process. The symbols we used in the rest of the paper are listed in Tab. I.
The overall LLM4AD framework is shown in Fig. 2. In our framework, humans provide instructions and evaluations, where the instructions $I$ and evaluations $F$ along with historical memories $H$, system messages $S$, and context information $C$ serve as inputs to LLMs. The memory module stores corresponding historical interactions between humans and vehicles $H$ for different users. Upon receiving these inputs, LLMs engage in reasoning and produce outputs, including generated Language Model Programs (LMPs) $P$ and reasoning thoughts $R$. The generated LMP is sent to the executor for execution in the environment, while the reasoning thoughts help LLMs generate more reasonable driving policies. It is important to note that this is a general concept, where specific implementations may vary across different applications. However, all applications follow the principles and framework of this concept.
II-A Human Instruction and Evaluation
Human instructions $I$ and evaluations $F$ are directly inputted to LLMs by humans in natural language. $I$ includes the humans’ desired needs for the autonomous agent, while human evaluations $F$ are the feedback on the effectiveness of these driving policies.
II-B System Message
The system message $S$ is a way to provide instructions or context to the LLM4AD system at the start of a conversation or task. System messages $S$ allow developers to clearly define the role, constraints, and objectives that LLM4AD systems should follow when engaging with users or working on particular types of tasks.
System messages $S$ in autonomous driving tasks act like a set of high-level guidelines or rules that frame how LLMs should operate and make decisions. These high-level system messages are crucial in defining the basic driving logic, including tasks such as task definition, adherence to traffic rules, description of decision states, and overall goals or metrics to optimize for. They serve as a foundational framework guiding the behavior and decision-making processes of autonomous vehicles on the road. Without a well-crafted system message, LLMs could make incorrect assumptions or unintended strategies.
II-C Situation Descriptor
A situation descriptor transfers the current driving context $C$ into textual descriptions. The situation descriptor follows a predefined structure and employs natural language to provide a detailed textual description of the current driving scenario. It aims to provide LLMs with situational awareness and a comprehensive representation of the current driving scenario, allowing them to make appropriate decisions in the current traffic situation. This could include descriptive statements like “You are positioned in the leftmost lane of a two-way highway” or “A vehicle is located 50 meters ahead of your current position.” The descriptor translates complex spatial and temporal relationships between different road users and ego vehicles intuitively into the natural language format, allowing context information to be reasoned by LLMs.
II-D History Memory and Memory Module
The memory module stores profiles from various users to enhance driving experiences for all users. Whenever a human user utilizes the LLM4AD system, the system logs the relevant historical interactions $H$ associated with that user. Subsequently, the historical data for the current human user is transmitted to LLMs as input. This particular history interaction $H$ serves as a reference point for the preferences of the current users, therefore guiding the system to improve user experience. After each trip, the interaction data will be updated in the corresponding profile in the memory module. The specifics of the interaction data vary for each system and will be elaborated in their respective sections.
II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human evaluation $F$) and generate the textual outputs (LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
II-F Generated Policy/Program
Inspired by the concept of “Code as Policy [liang_code_2023],” one of the key outputs from LLMs is the generated LMPs $P$ consisting of executable codes. These codes are used to affect the driving behavior of the ego agent within the environment, and they not only have the ability to generalize to new natural language commands but can also provide precise numerical values such as velocities based on vague linguistic descriptions (e.g., “hurry,” “turn left”) depending on the driving context.
II-G Output Thought
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is “hurry,” I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated programs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.
II-H Executor
The executor acts as the bridge between LLMs’ textual outputs and the current autonomous driving policy. It takes the generated LMPs $P$ from LLMs and executes it in the corresponding environment. This allows the codes to interact with the current status of the ego vehicle and enables the generated programs to deploy their intended driving behaviors in the real or simulated environment. In our scope, different autonomous driving systems will have some differences in how those codes are executed by their respective executors.
III Benchmarking LLM4AD in Simulation
This section introduces two comprehensive benchmarks for evaluating the instruction-following capabilities and performance of LLM-based agents in autonomous driving: the LaMPilot-Bench dataset on the HighwayEnv simulator and the CARLA Leaderboard 1.0 [dosovitskiy_carla_2017].
III-A LaMPilot-Bench
III-A1 Dataset
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. II. For each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed.
TABLE II: Main statistics of the LaMPilot-Bench dataset.
Statistic Value
Total scenarios 4,900 (100%)
Distance-related 1,200 (24.5%)
Speed-related 1,200 (24.5%)
Pulling over 200 (4.1%)
Routing 1,500 (30.2%)
Lane changing 400 (8.2%)
Overtaking 400 (8.2%)
Highway 3,400 (69.4%)
Intersection 1,500 (30.6%)
Average instruction length 7.6 words
Maximum instruction length 14 words
Minimum instruction length 2 words
Figure 3: Distribution of instructions in the LaMPilot-Bench dataset.
The LaMPilot-Bench simulator builds upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
III-A2 Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) measures the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 0), the TTC ($\tau_{i}$) with vehicle $i$ is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
$\tau_{i}=-\frac{(\mathbf{p}_{0}-\mathbf{p}_{i})\cdot(\mathbf{v}_{0}-\mathbf{v}_{i})}{\|\mathbf{v}_{0}-\mathbf{v}_{i}\|^{2}}$ (1)
The minimum positive TTC value $\tau_{\min}$ among all $n$ vehicles over all time steps $t$ (up to task completion time $T$) represents the nearest potential collision:
$\tau_{\min}=\min_{1\leq i\leq n,1\leq t\leq T}\tau^{t}_{i}$ (2)
TTC scores are empirically assigned based on a 2-second safety margin, with values above 2 seconds considered safe and given a score of 100:
$\operatorname{TTC}=\begin{cases}100&\text{if $\tau_{\min}>2$}\\ \max(0,100-\frac{1}{\tau_{\min}})&\text{if $0<\tau_{\min}\leq 2$}\\ 0&\text{if $\tau_{\min}\leq 0$}\end{cases}$ (3)
Speed variance (SV) is another safety metric. The ego vehicle’s speed standard deviation $\sigma_{0}$ is calculated as:
$\sigma_{0}=\sqrt{\frac{\sum_{t=1}^{T}{(\|\mathbf{v}^{t}_{0}\|-\mu_{0})^{2}}}{T}},$ (4)
where $\mu_{0}$ is the mean speed:
$\mu_{0}=\frac{\sum_{t=1}^{T}{\|\mathbf{v}^{t}_{0}\|}}{T}.$ (5)
The SV score is defined as:
$\operatorname{SV}=\max(0,100\cdot(1-{\sigma_{0}}/{\sigma_{\operatorname{safe}}})),$ (6)
where $\sigma_{\operatorname{safe}}$ is the maximum safe speed deviation, determined empirically. The time efficiency (TE) score evaluates the policy’s ability to complete the task within a predefined time limit $T_{\operatorname{limit}}$:
$\operatorname{TE}=\max(0,100\cdot(1-{T}/{T_{\operatorname{limit}}}))$ (7)
A TE score closer to 100 indicates a more efficient policy. A task is considered successfully completed when the agent achieves the specified objectives while maintaining safety (i.e., avoiding collisions) and efficiency (i.e., finishing within the allotted time). The final score aggregates all individual metrics, weighted according to their importance:
$\operatorname{Score}=W_{\operatorname{TTC}}\cdot\operatorname{TTC}+W_{\operatorname{SV}}\cdot\operatorname{SV}+W_{\operatorname{TE}}\cdot\operatorname{TE}$ (8)
III-B CARLA Leaderboard 1.0 Benchmark
In addition to LaMPilot-Bench and HighwayEnv, we also utilize the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 training, 26 testing) totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, we modify the Leaderboard setup to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, we craft the driving context information using a language generator that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
  • Route Completion (RC): measures the percentage of the route distance completed by the agent.
  • Infraction Penalty (IP): tracks various infractions (e.g., collisions, running red lights) committed by the agent, aggregated as a geometric series starting from an ideal base score of 1.0 that is reduced with each infraction.
  • Driving Score (DS): the product of RC and IP, serving as the principal evaluation metric.
III-C Baselines and Results
We establish several baselines to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. III for LaMPilot-Bench and Tab. IV for the CARLA Leaderboard 1.0.
TABLE III: Comparison of baselines on LaMPilot-Bench under different learning settings.
Method Learning Collision ($\downarrow$) Completion ()($\uparrow$)( ↑ ) TTC Score ()($\uparrow$)( ↑ ) SV Score ()($\uparrow$)( ↑ ) TE Score ()($\uparrow$)( ↑ ) Driving Score ()($\uparrow$)( ↑ )
IDM N/A 0.0% 20.4% 92.8 80.9 71.0 17.3
MOBIL 0.0% 15.3% 82.8 85.9 46.7 11.7
Human Driver 0.0% 98.0% 93.4 86.3 81.3 84.6
Llama 2 0-Shot 0.0% 20.4% 92.8 81.4 68.9 17.3
PaLM 2 3.1% 35.7% 78.6 76.5 78.8 12.8
ChatGPT 1.0% 40.8% 83.9 75.5 74.0 27.8
GPT-4 4.1% 72.4% 61.3 70.5 74.0 28.5
GPT-4 Turbo 3.1% 74.5% 73.2 70.1 73.9 39.1
Llama 2 3-Shot 1.0% (+1.0) 63.3% (+42.9) 68.3 (-24.5) 71.4 (-10.0) 73.8 (+4.9) 39.6 (+22.3)
PaLM 2 3.1% (0.0) 71.4% (+35.7) 73.4 (-5.2) 69.7 (-6.8) 72.0 (-6.8) 36.6 (+23.8)
ChatGPT 4.1% (+3.1) 83.7% (+42.9) 70.9 (-13.0) 73.7 (-1.8) 77.7 (+3.7) 41.2 (+13.4)
GPT-4 1.0% (-3.1) 84.7% (+12.3) 69.4 (+8.1) 72.0 (+1.5) 76.7 (+2.7) 55.9 (+27.4)
GPT-4 Turbo 1.0% (-2.1) 80.6% (+6.1) 69.1 (-4.1) 71.5 (+1.4) 74.9 (+1.0) 52.6 (+13.5)
Llama 2 Human Feedback 0.9% (-0.1) 32.7% (-30.6) 83.9 (+15.6) 74.6 (+3.2) 74.6 (+0.8) 21.4 (-18.2)
PaLM 2 0.0% (-3.1) 45.5% (-25.9) 81.1 (+7.7) 74.4 (+4.7) 77.6 (+5.6) 35.6 (-1.0)
ChatGPT 0.9% (-3.2) 80.9% (-2.8) 70.2 (-0.7) 71.7 (-2.0) 77.5 (-0.2) 54.2 (+13.0)
GPT-4 0.9% (-0.1) 92.7% (+8.0) 67.6 (-1.8) 72.2 (+0.2) 76.8 (+0.1) 64.0 (+8.1)
GPT-4 Turbo 0.9% (-0.1) 87.3% (+6.7) 74.1 (+5.0) 73.4 (+1.9) 75.5 (+0.6) 60.5 (+7.9)
TABLE IV: Driving performance of expert drivers on CARLA leaderboard 1.0 testing routes. $\uparrow$: Higher values are better.
Method Learning Navigation DS ()($\uparrow$)( ↑ ) RC ()($\uparrow$)( ↑ ) IP ()($\uparrow$)( ↑ )
Roach Expert [zhang_end--end_2021] Reinforcement Learning Waypoint 63.4 100.0 0.63
TCP Expert [wu_trajectory-guided_2022] Reinforcement Learning Waypoint 68.1 95.8 0.71
LLM Planner 0-Shot CoT Instruction 0.00 48.7 0.00
LLM Planner 3-Shot CoT Instruction 17.3 65.3 0.31
LLM Planner Human-Feedback CoT Instruction 65.4 95.1 0.68
Heuristic baselines
Rule-based methods such as the Intelligent Driver Model (IDM)[treiber_congested_2000] and Minimizing Overall Braking Induced by Lane Changes (MOBIL)[kesting_general_2007]. Both methods achieve a zero collision rate in LaMPilot-Bench, as shown in Tab. III. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
Zero-shot and few-shot baselines
LLMs are prompted with task instructions and API documentation, with or without in-context examples. For the few-shot setting, human programmers proficient in Python create in-context examples. They are given API documentation and allowed to write and test their code on the training routes. The same examples are used across all test routes. In the zero-shot setting, GPT models [openai_gpt-4_2023] and PaLM 2 [anil_palm_2023], given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion, as shown in Tab. III. When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2 [touvron_llama_2023]’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting.
Human feedback baselines
LLMs are integrated with a human-in-the-loop learning approach, leveraging human feedback to refine generated policies. After executing a generated policy code ($P$), the human evaluators provide natural language feedback ($F$), which is fed back into the LLM along with $P$. This feedback loop enables continuous learning. If the feedback is positive (i.e., the human is satisfied with the execution), the code ($P$) is committed to the database for future retrieval and reuse. Otherwise, the feedback serves as guidance for iterative improvement. The integration of GPT models with human-in-the-loop learning further enhances their performance in driving tasks. Notably, GPT-4 achieves the highest driving score of $64.0$, as shown in Tab. III. This improvement highlights the great potential of LLMs in following instructions in the driving context when combined with human feedback.
Human performance baseline
A licensed human driver controlling the vehicle in the simulation using keyboard inputs. This baseline provides a reference for human-level performance.
Comparison with other methods
We compare the human-feedback baseline with reinforcement learning experts that have access to privileged information, such as Roach [zhang_end--end_2021] and TCP [wu_trajectory-guided_2022] in the CARLA leaderboard [ma_learning_2024], as shown in Tab. IV. The comparison aims to provide a reference to previous works. Based on these results, we observe the following:
  • Without few-shot examples, the out-of-the-box LLMs struggle with the precise reasoning required for complex closed-loop driving in CARLA.
  • The three-shot baseline also falls short, resulting in significantly lower scores according to CARLA’s standard metrics.
  • However, the LLM with 50 code snippets learned from human feedback performs on par with the Roach Expert reinforcement learning baseline.
The experiments demonstrate that off-the-shelf LLMs can generate code policies for driving tasks based on human instructions, outperforming rule-based methods in task completion. Few-shot learning with exemplar code further improves the performance of LLM-based agents, and the integration of human feedback enables iterative refinement of generated policies, leading to the highest driving scores among the evaluated methods. However, the notable collision rates in both benchmarks indicate the need for further research to fully capture the complexities and safety requirements of real-world driving scenarios. Most failures stem from the trade-off made for the LLM low-level planning frequency, as illustrated in the example below:
# Context Info:
# There is a vehicle at a distance of 6.8 m.
# It is moving at a speed of 0.0 m/s.
# It is at an angle of 1.9 degrees.
# The angle between the headings is 0.0 degrees.
# My current speed is 0.0 m/s.
def start_driving_with_front_vehicle_check():
(is_front_vehicle,
distance_to_front_vehicle,
speed_of_front_vehicle)
= check_front_vehicle()
if (is_front_vehicle
and distance_to_front_vehicle < 5
and speed_of_front_vehicle == 0.0):
yield stop()
else:
speed_limit = check_speed_limit()
yield proceed(speed_limit)
[…]
In this case, the vehicle could collide with a front vehicle if it starts to move. This issue arises because the LLM reasoning is based on the current context information and does not account for all potential future actions of surrounding agents. Due to the latency of LLM reasoning, it is not feasible to employ LLMs to reason at every single timestamp, which is a challenge to be addressed in future research.
III-D Summary
The LaMPilot-Bench dataset and the CARLA Leaderboard 1 benchmark provide a comprehensive evaluation framework for assessing the instruction-following capabilities and performance of LLM-based agents in AD. The experiments highlight the potential of LLMs in enhancing the decision-making capabilities of autonomous vehicles when combined with human feedback and domain-specific APIs. As LLMs continue to evolve and integrate with more modalities, we envision their significant contribution to the development of safe, efficient, and user-friendly autonomous vehicles. Future research should focus on addressing the challenges of real-time decision-making, reducing collision rates, and incorporating multi-modal inputs to further enhance the performance and robustness of LLM-based autonomous driving systems.
Figure 4: The proposed framework for personalized decision-making with an on-cloud large language model in real-vehicle experiments.
IV Real-Vehicle Experiments: On-Cloud LLM for Personalized Decision-Making
To further assess the effectiveness of the LLM4AD system and validate its applicability in real-world scenarios, we integrate LLMs into practical autonomous driving systems, introducing a framework called Talk2Drive. Similar to the concept in Sec. II, our system has outstanding command understanding and reasoning performance. It also can provide a personalized driving experience by leveraging records of previous interactions. The demo for our real-vehicle experiment is publicly available [^1] Real-vehicle experiment demo video with on-cloud LLM: https://youtu.be/4BWsfPaq1Ro.
IV-A Talk2Drive Framework
This section proposes Talk2Drive (see Fig. 4), an innovative approach to leveraging LLMs to enhance command interpretation and enable personalized decision-making in autonomous vehicles. It integrates cloud-based LLMs to enable personalized understanding and translation of human commands into executable control sequences with real-time vehicle dynamic inputs. This section begins with a problem statement and then articulates the distinct roles of each cloud-side and vehicle-side operation. The flowchart of the Talk2Drive system is shown in Fig. 5.
IV-A1 Problem Statement
In Talk2Drive, we follow the concept we proposed in Sec. II. The model aims to translate verbal commands into executable control sequences for the vehicle. The cloud-based LLM acts as a translating function $f:\boldsymbol{I}\rightarrow\boldsymbol{P}$ that generates corresponding LMPs as the policy ($\boldsymbol{P}$) for maneuvers.
For personalization, we need human feedback on execution $F(I,P)$ to evaluate if the generated policy addresses the preferences of the driver or the passengers. The extra memory module (see IV-A5) includes instructions, LMPs, and their evaluation will allow to guide the LLM to learn preferences. Therefore, there are two stages in the Talk2Drive workflow:
[…]
IV-A2 Command Translation and Contextual Data Integration
The initial step in the Talk2Drive framework involves directly receiving arbitrary verbal commands from humans. Utilizing cutting-edge voice recognition technology, specifically the open-source API Whisper [radford2023robust], these verbal commands are accurately captured and then translated into textual instructions ($I$). An instance for $I$ is:
$I\to\text{Could you drive more conservatively?}$? (10)
Simultaneously, LLMs access additional cloud-based real-time environment data, including weather updates, traffic conditions, and local traffic rules. In our product, LLMs can be informed by the weather information through Openweather API [openweather], the map information (such as road type and speed limits) through OpenStreetMap API [OpenStreetMap], and traffic information through TomTom API [tomtom]. We use a predefined structured language generator for both $C$ and $S$, and our system will supply the appropriate values (the red contents) to the generator based on the context information, as shown below:
$C\to$ (11)
$\left\{\begin{array}[]{l}\text{A vehicle in front of you is running at {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}38.0} km/h.}\\ \text{Your current speed is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}40.0} km/h.}\\ \text{The speed limit is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}60.0} km/h.}\\ \text{The weather is {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}sunny}.}\\ \vdots\end{array}\right.$ (17)
IV-A3 Processing and Reasoning with LLMs
The LLMs are prompted using in-context learning coupled with chain-of-thought prompting. Specifically, the chain of thought’s prompt in Talk2Drive consists of triples: $\{query,thought,output\}$. An example can be found in Eq. 18.
Once a command is translated into texts, it is uploaded to LLMs hosted on the cloud. In our experiments, we utilize GPT-4 [openai_gpt-4_2023] as our LLM and access it through ChatGPT API. The LLMs engage in a reasoning process to interpret the instruction ($I$), the contextual data ($C$), the system messages ($S$), and the history data ($H$). A simplified example of $S$ is presented below:
$S\to$ (18)
$\left\{\begin{array}[]{l}\text{You are an autonomous vehicle withAdaptive cruise}\\ \text{control (ACC) and Lane Keeping Assist (LKA) always}\\ \text{enabled.}\\ \text{You are using Pure Pursuit Controller to do the waypoint}\\ \text{following.}\\ \vdots\\ \text{Here are some examples of how you need to react.}\\ \text{Query: You drive too aggressively.}\\ \text{Thought: The drivers think I drive too fast which looks}\\ \text{aggressive and the drivers do not ask me to change lanes,}\\ \text{so I need to slow down my speed.}\\ \text{Action: ...}\\ \vdots\end{array}\right.$ (32)
Additionally, the LLMs also access the memory module, a repository of historical interactions ($H$), to consider the human’s past behaviors and preferences. More details about the memory module are in Sec. IV-A5.
IV-A4 Actionable Code Generation
Same as in Sec. II, the LLM is utilized to generate corresponding LMPs ($P$) based on this interpretation. These LMPs include complex driving behaviors and parameters that need to be adjusted in the vehicle’s high-level controllers. Specifically, the LMPs adjust control parameters like the look-ahead distance and look-ahead ratio to optimize pure pursuit [pure_pursuit] performance in Talk2Drive. Additionally, LMPs also modify the target velocity of the vehicle to meet humans’ commands. These LMPs take the form of ROS topic commands, directing the autonomous driving system based on Autoware [autoware] to modify its trajectory following configuration.
One simple example codes are shown as follows:
$P\to$ (33)
\$ timeout 1s rostopic pub /vehicle/engage
std_msgs/Bool "data: true"
\$ rostopic pub /autoware_config_msgs
/ConfigWaypointFollower
"{\"param_flag\": 1, \"velocity\": 40,
\"lookahead_distance\": 12,
\"lookahead_ratio\": 2.0}"
In the meantime, LLMs also generate thoughts $R$ to explain why they make a particular decision. According to  [wei_chain--thought_2022], even though the thoughts $R$ are not executed in our systems, they help LLMs make more reasonable and feasible decisions. An example of the thought $R$ is shown below:
$R\to$ (34)
$\left\{\begin{array}[]{l}\text{The driver wants me to drive more conservatively but}\\ \text{does not have any specific lane preference. Therefore, I}\\ \text{should reduce my speed and maintain my current lane.}\\ \end{array}\right.$ (38)
The generated LMPs ($P$) are then sent back from the cloud to the vehicle’s ECU, where they are executed. Additionally, we also set two kinds of safety checks for the generated LMPs. The first check involves verifying the format of the generated code. If the code does not conform to a valid format, our system will not execute or respond to the code. The second safety check is parameter verification, which assesses the feasibility and safety of the parameters within the context of the given policy, preventing the execution of potentially risky policies. For example, if the generated code instructs our agent to move to the left lane while the ego agent is already in the leftmost lane, our system will identify this abnormality and prevent the execution.
IV-A5 Memory Module and Personalization
As described in Sec. II, the memory module is used to store the historical interactions ($H$) between humans and vehicles to provide a personalized driving experience. One important difference to note about this specific system, compared to the general framework, is that within this system, human evaluations $F$ are archived within the historical memory, and the history memory $H$ includes three key linked components: the past commands ($I$) issued by the human, the corresponding LMP ($P$) generated by the LLMs in response, and the humans’ subjective evaluations or feedback ($F$) on the system’s performance for each command. Each interaction between the human and the vehicle is recorded and saved into a memory module in a text format within the ECU. This historical data in the memory module is updated after every trip. The example contents in the memory module are shown in Eq. 39:
$H\to$ (39)
$\left\{\begin{array}[]{l}\text{Apart from the requirements I provided before. Here I}\\ \text{will provide you with the history dialogues between the}\\ \text{driver and the vehicle. You will need to learn what}\\ \text{drivers' wants and needs are.}\\ \text{For example, if the history driver's command is ``You}\\ \text{drive too conservatively.''}\\ \text{The history output action is : ...}\\ \text{After the trip, the driver's feedback is ``A little bit too}\\ \text{fast for me.''}\\ \text{Then the next time your action should adjust according}\\ \text{to the driver's feedback, which is ...}\\ \vdots\\ \text{The history command, action, and driver's feedback are:}\\ \text{Command: I'm on my way to the urgent care.}\\ \text{Action: ...}\\ \text{Evaluation: A little bit too fast.}\\ \vdots\end{array}\right.$ (57)
Figure 5: The flowchart of Talk2Drive. After the speech recognition module detects the keyword “command,” the inputs ($I,C,S,H$) are sent to the LLM. Then, the LLM generates corresponding LMPs to be executed by the ECU. If the speech recognition module detects the keyword “evaluate,” the system receives human feedback ($F$), and both $F$ and its corresponding $I$ and $P$ are updated in the memory module.
Given the adaptive nature of LLMs, if users respond differently to similar commands, the LLMs will prioritize the most recent response as a reference point for their current decision-making process. When a command from the users is issued, the LLMs access the memory module and take stored information ($H$) as part of the input prompts for the decision-making process. Additionally, each user has their own profile in the memory module, ensuring that our framework can deliver personalized driving strategies for everyone.
Figure 6: Experiment scenarios: car following on a three-lane highway, unprotected left turn at an intersection, and event parking at a parking lot.
Figure 7: Setup of the real-world autonomous vehicle in the experiment.
IV-B Autonomous Driving System
As shown in Fig. 7, we use an autonomous vehicle to conduct real-world experiments, which is a drive-by-wire enabled 2019 Lexus RX450h with perception sensors, localization module, and communication module. We deploy the open-source autonomous driving software Auoware.AI [autoware] with ROS Melodic in Ubuntu 18.04. We employ 3D-NDT [3Dndt] for mapping and localization, and we utilize pure pursuit [pure_pursuit] for trajectory following. Network communication is enabled by the Cradlepoint IBR900 Series Router, which is equipped with an AT&T sim card with a 4G-LTE cellular connection.
IV-C Experiment Setup
IV-C1 Scenarios Setup and Participants
Our experimental trials include three different scenarios: highway, intersection, and parking lot.  [^2] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. The field experiments for the highway and intersection scenarios are conducted at a proving ground in Columbus, IN, USA, to validate the efficacy of the proposed Talk2Drive framework. The highway scenario involves a three-lane highway, while the intersection contains a two-way junction. Additionally, the parking lot scenario is evaluated at the North Stadium Parking Lot in West Lafayette, IN, USA. The overview visualization and statistics of the test tracks are shown in Fig. 6, while the visualization for the experiments is shown in Fig. 8. In our study, we recruit a diverse range of drivers, including individuals of different genders (61.4% male, 28.6% female), ages (mean=26.71, std=4.11), and driving experiences(mean=6.79, std=5.08). All participants have valid driving licenses.
IV-C2 Input Instructions
In the field of linguistics, giving instructions can be categorized as making a request, which falls under the division of directives, one of the five speech act types according to the general speech act classification theory [yule2022study]. More specifically, the scale of directness in requests can be characterized by three distinct strategies [shoshana1989cross]:
Direct Strategies: In this scenario, the human explicitly states the desired action, such as “Increase the speed of the vehicle.” This is usually in the form of imperative sentences.
Conventionally Indirect Strategies: This scenario involves phrasing requests in a manner that is socially and culturally acknowledged as polite and/or standard. An example would be “Could you please speed up a little bit?”
Non-Conventionally Indirect Strategies: Requests in this scenario are more implicit and rely on contextual understanding. Within this category, hints can be further divided into strong and mild hints. In the context of requesting speed changes in an autonomous vehicle, strong hints may include explicit comments on the current speed, such as “You are driving too aggressively.” Conversely, mild hints might be more oblique references to time or urgency, for example, “I hope we will not be late for the meeting.”
Given that conventionally indirect strategies primarily modify the politeness level of command of direct strategies, we instead divide the non-conventionally indirect requests into two categories based on the strength of the hints. We define our levels of directness in Tab. V as our way to classify varying degrees of implicitness in the commands. To gather data, we request our test drivers to generate commands based on their normal speaking preferences and subsequently categorize their commands into our defined levels. Examples of the commands generated are also presented in Tab. V.
Figure 8: The experiment visualization: In the upper left corner is the in-cabin view, while the lower left corner displays the exterior view. The upper right corner shows the console, and the lower right corner presents the lidar map.
TABLE V: Level of Command Directness and Examples
Command Level Linguistic Category Example Commands
Level I Direct and conventionally indirect commands Drive as fast as you can.
Can you drive faster?
You should drive slower.
Level II Non-conventionally indirect commands with strong hints You are driving too conservatively.
You are driving a little bit slow now.
I do not think we should drive this slow.
Level III Non-conventionally indirect commands with mild hints I am going to be late for my work.
I feel a bit motion-sick right now.
I’m really in a hurry now.
TABLE VI: Driving Performance Validation
[…]
Time to
Collision($s$$\uparrow$
Speed
Variance($m^{2}/s^{2}$$\downarrow$
Mean Absolute
Acceleration($m/s^{2}$$\downarrow$
Mean Absolute
Jerk($m/s^{3}$$\downarrow$
LLM
Latency($s$$\downarrow$
Highway Overtake I 2.88 5.02 0.24 2.64 1.68 85.05
II 1.94 4.05 0.24 2.81 1.87 86.12
III 3.07 1.26 0.18 2.64 1.86 91.12
Baseline 3.26 2.91 0.35 2.83 - 86.00
Following I 6.52 0.94 0.15 2.35 1.61 87.10
II 7.84 1.11 0.15 2.38 1.64 86.23
III 6.78 1.37 0.09 2.31 1.64 86.26
Baseline 4.02 0.78 0.22 2.50 - 86.00
Right Lane I 8.77 1.69 0.17 2.39 1.32 90.88
II 4.54 1.18 0.15 2.44 1.83 91.51
III 7.29 0.23 0.13 2.61 1.22 92.18
Baseline 4.70 7.39 0.22 2.77 - 86.00
Intersection Not Yield I 0.89 0.29 0.26 2.28 1.65 59.67
II 0.89 0.22 0.28 2.55 1.78 59.60
III 1.04 0.21 0.26 2.32 1.52 60.32
Baseline 1.14 0.46 0.46 2.34 - 56.00
Yield I - 0.29 0.52 2.27 1.47 91.53
II - 0.22 0.82 2.54 1.43 89.50
III - 0.21 0.48 2.28 1.38 91.92
Baseline - 1.67 0.90 2.32 - 86.00
IV-C3 Evaluation Metrics
Our evaluation framework for autonomous vehicles includes driving performance, time efficiency, and personalization. We analyze Talk2Drive’s driving performance in terms of safety, using Time to Collision ($\tau$) and speed variance ($\sigma^{2}$), and comfort, using mean absolute acceleration ($|\bar{A}|$) and mean absolute jerk ($|\bar{J}|$). Time efficiency is measured by the LLM latency ($L$), while human satisfaction with personalization is assessed using the takeover rate ($R$).
Driving Performance The overall driving performance of Talk2Drive is reflected by a driving score ($S$), which is a weighted sum of four sub-scores: Time to Collision score ($S_{\tau}$), speed variance ($S_{\sigma}$), mean absolute acceleration ($S_{|\bar{A}|}$), and mean absolute jerk ($S_{|\bar{J}|}$).
$S=\sum{w_{k}\cdot S_{k}}$ (58)
For Time to Collision $\tau$, the critical threshold $\tau_{c}$ is based on human reaction time to take over and brake. A TTC greater than 1.5 seconds is often considered safe as the driver will have enough time to react to avoid rear-end collisions [van1994time]. Therefore, it can be considered as a hard threshold where any value greater than $\tau_{c}$ has a score of zero.
$S_{\tau}=\begin{cases}0,&\text{if }\tau_{min}<\tau_{c}\\ 100,&\text{if }\tau_{min}\geq\tau_{c}\end{cases}$ (59)
where $\tau_{c}$ is the critical threshold. We chose the value of 1.5 based on existing research results on Time to Collision value [van1994time]. For other metrics like speed variance, acceleration, and jerk, the thresholds are related to individual perception and would vary in different driving scenarios. Therefore, we define the score to quantify the corresponding performance, with a higher score indicating better performance relative to the baseline value. For example, the sub-score for mean absolute jerk $|\bar{J}|$ can be denoted as:
$S_{|\bar{J}|}=100-\gamma\cdot\frac{|\bar{J}|}{|\bar{J}|_{\text{baseline}}}$ (60)
where $\gamma$ is the sensitivity factor which is set empirically. The sub-scores for mean absolute acceleration and speed variance are obtained in the same way.
LLMs’ Latency for Time Efficiency Latency $L$ measures the response time of the LLM, which is important for time-sensitive vehicle applications. To measure latency, we calculate the time difference between the moment a command is sent to the LLMs and when the LLMs return the LMPs.
Takeover Rate for Personalization The frequency of manual interventions indicates the model’s ability to adapt to personalized preferences from different humans [wang2022gaussian].
IV-D Experiment Results
In the experiments, all participants are required to issue verbal commands to the systems, which will be either baselines or Talk2Drive systems chosen randomly. To avoid bias, the participants will be unaware of which system is in use. They can decide when to take over the driving system, and the takeover rate will be recorded. Takeovers in this context are noted when humans find the current autonomous system unsatisfactory. Additionally, all driving data will be logged, and the ping latency for our experiments is at 200$\sim$400 ms.
IV-D1 The Validation of Driving Performance of Talk2Drive
To validate the driving performance of the Talk2Drive system, we conduct comprehensive experiments to compare the driving performances between the baseline and the Talk2Drive through the metrics defined in Sec. IV-C3. As discussed earlier, the baseline values are from the average data of human drivers. The average driving performance for each scenario within each command directness is shown in Tab. VI. Note that for each expected driving behavior, the driving pattern is distinct. Therefore, a baseline is created for every driving behavior separately.
The driving score is the weighted sum of all the safety and comfort metrics, reflecting the overall driving performance. The driving scores are calculated and recorded in Tab. VI, with green texts highlighting scores higher than the baseline of the same driving behavior and red texts showing the lower scores. As shown in Tab. VI, regardless of the command’s directness or the human’s behavior, nearly all driving scores are better than the baseline, showing that our framework consistently provides relatively safe and comfortable driving experiences for humans in all scenarios.
The latency of the internet remains mostly consistent (around 300ms) under the same network conditions. The empirical results from our tests reveal that the latency for the most responsive LLM falls between approximately 1.2 to 1.8 seconds. This finding suggests considerable potential for integration into certain real-time applications where such response times are within acceptable limits. Specifically, LLMs demonstrate strong suitability for making high-level decisions or modifying driving behaviors in response to human commands. These scenarios require timely responses but do not typically demand extremely low latency.
IV-D2 The Improvement on Personalization Performance
This section explores the impact of integrating the Talk2Drive framework into autonomous driving systems to enhance personalization based on the LLM model GPT-4 [openai_gpt-4_2023]. One of the central focuses of our system is its ability to offer drivers a personalized driving experience. Through Talk2Drive, individuals can express their preferences or feelings with varying degrees of directness, prompting the system to make corresponding adjustments. The commands collected are divided into three levels of directness, as explained in Tab. V.
For each command directness within each expected driving behavior, the system’s responses to various responses given by human participants are collected and calculated into the defined metrics. Here, we utilize the takeover rate as our primary metric, and we assess the takeover rate both pre- and post-integration of Talk2Drive. We incorporate a rule-based system with a keyword-trigger logic as the baseline for comparison. For instance, in the highway scenario, commands like “accelerate,” “left,” and “right” prompt the system to execute relevant actions such as speeding up or changing lanes accordingly. Results are gathered in diverse driving scenarios on human drivers with diverse driving styles.
TABLE VII: Takeover Rate Improvement through Personalization.
Driving
Scenario
Command
Directness
Baseline Talk2Drive Reduction
Highway I 0.33 0.07 78.8%
II 0.63 0.20 68.3%
III 0.77 0.31 59.7%
Intersection I 0.33 0.11 66.7%
II 0.71 0.29 59.2%
III 0.48 0.21 56.3%
Parking I 0.07 0 100%
II 0.20 0 100%
III 0.67 0.24 64.2%
As shown in Tab. VII, the takeover rate in all driving scenarios decreases significantly, ranging from a 56.3% reduction to the complete elimination of takeover cases, demonstrating the Talk2Drive system’s ability to personalize the driving experience based on human preference. Additionally, one important finding from our results is that for all levels of command directness, our Talk2Drive system shows significant improvements compared to the baseline systems.
IV-D3 The Effectiveness of the Memory Module
To further investigate the performance of our memory module, we also conduct experiments comparing personalization across two settings: conditions without the memory module, and conditions with the memory module. These experiments take place in the parking lot, where the vehicle follows a pre-defined trajectory and adjusts speed based on human inputs. Humans take over in this context when they find the adjusted speed unsatisfactory. As demonstrated in Tab. VIII, there are significant reductions in takeover rate with Talk2Drive, regardless of the driver’s aggression or conservatism levels.
The results in Tab. VIII reveal that the inclusion of the memory module leads to a marked reduction in the takeover rate. The utilization of Talk2Drive framework without the memory module brings takeover rates between 0.14 and 0.29. When implementing Talk2Drive framework with the memory module, the takeover rate can be further decreased to only 0.07. Compared to their personalization performance, Talk2Drive with a memory module can reduce takeover rates by up to 65.2% compared to those without the memory module, which illustrates the benefits of a history-recording module in achieving a more personalized driving experience.
TABLE VIII: Effectiveness of Memory Modules in Influencing Takeover Rates
Driver Without Memory Module With Memory Module
A 0.14 0.07 (50.0$\downarrow$)
B 0.23 0.08 (65.2$\downarrow$)
C 0.29 0.18 (37.9$\downarrow$)
V Real-Vehicle Experiments: On-Board VLM for Personalized Motion Control
In this section, we present our on-board VLM for personalized motion control in autonomous driving, designed to accommodate individual driving styles. Our approach leverages a compact 8B-parameter VLM, fine-tuned from Qwen-VL [bai2023qwenvlversatilevisionlanguagemodel], which processes both visual information (including weather conditions, road types, and traffic conditions) and verbal commands to generate personalized control strategies for each user. The reduced scale of this VLM enables edge deployment while maintaining command interpretation and reasoning capabilities, allowing the system to effectively understand and respond to implicit human instructions. The overall framework is shown in Fig. 9.
Figure 9: The proposed framework for personalized motion control with an on-board vision language model in real-vehicle experiments.
V-A Problem Statement
We adopt the traditional module-based autonomous driving framework that includes the full pipeline from perception to motion control, and our focus is specifically on enhancing the decision-making process at the motion control level, adapting the control performance to accommodate different human driving styles. The goal of the proposed system is to translate both verbal commands $\boldsymbol{I}$ and visual inputs $\boldsymbol{V}$ into executable control sequences for the motion control process. The onboard VLM acts as a translation mechanism $f:(\boldsymbol{I},\boldsymbol{V})\rightarrow\boldsymbol{P}$ that generates a policy $\boldsymbol{P}$, which is then fed into predefined maneuver programs.
Additionally, system messages $\boldsymbol{S}$ are sent to our VLM to specify both the tasks and adjustment strategies. In our setting, similar to the Sec. IV, $\boldsymbol{S}$ is generated through a predefined language generator.
Simultaneously, to further enhance personalization, we implement a RAG-based memory module to build a database storing historical human-vehicle interactions. Whenever a human activates this system, only relevant historical scenarios $\boldsymbol{H}$ are retrieved and provided to the VLM as reference. After each trip, users can provide feedback $\boldsymbol{F}$ on the generated control policy $\boldsymbol{P}$ for the current situations (including instructions $\boldsymbol{I}$ and visual input $\boldsymbol{V}$), which helps refine the VLM’s reasoning process. Subsequently, the instructions $\boldsymbol{I}$, scene description $\boldsymbol{D}$ (generated by on-board VLM), policy $\boldsymbol{P}$, and feedback $\boldsymbol{F}$ are packaged as a historical vector-based data entry and stored in the RAG database. Therefore, there are three procedures in our system:
[…]
V-B Our Framework
V-B1 System Input
As illustrated in Fig. 9, our fine-tuned on-board VLM processes four critical inputs for reasoning. The primary inputs consist of visual data $\boldsymbol{V}$ from the onboard camera and natural language commands $\boldsymbol{I}$ which are converted from verbal human instructions $\boldsymbol{I}$ using the open-source local speech recognition model “Whisper [radford2022robustspeechrecognitionlargescale].” Notably, our system leverages contextual and environmental information captured in the visual inputs $\boldsymbol{V}$, including weather conditions, traffic conditions, and road characteristics. For instance, the system automatically adopts a more conservative driving policy during nighttime operations or adverse weather conditions.
Additionally, a pre-defined system message generator is employed to produce a customized system message $\boldsymbol{S}$, which is then simultaneously sent to the VLM.
Furthermore, the VLM incorporates relevant interaction history $\boldsymbol{H}$ extracted from our RAG-based memory module as contextual input, which includes previous human instructions $\boldsymbol{I}$, scene descriptions $\boldsymbol{D}$, executed actions $\boldsymbol{A}$, and user feedback $\boldsymbol{F}$. Note that the scene descriptions $\boldsymbol{D}$ are generated by the on-board VLM to describe the image $\boldsymbol{V}$. This historical context enables the VLM to generate more appropriate responses by considering past interactions and human feedback. A detailed discussion of how our memory module works will be presented in subsection V-B3.
V-B2 Reasoning and Action Generation
In our approach, reasoning within the VLM framework enables the interpretation of diverse driving scenarios and user instructions to generate actionable outputs. Traditional controllers in motion control typically rely on a default set of parameters; however, following the approach in [sha2023languagempclargelanguagemodels], our VLM will generate two distinct action matrices to separately manage the PID controller for longitudinal movement and the MPC for lateral movement. These matrices translate the model’s understanding of the environment and user preferences into precise control actions, guiding the autonomous vehicle’s behavior. Specifically, they are used by the controllers to generate acceleration $a$ and steering angle $\delta_{f}$, which are executed by the vehicle’s ECU. The ECU then sends low-level control signals to the drive-by-wire system developed by AutonomousStuff [autostuff], enabling smooth and responsive vehicle operation. The general process of this subsection can be shown below:
Output Action $P:=\begin{bmatrix}K_{p}&K_{i}&K_{d}\\ W_{l}&W_{h}&W_{s}\end{bmatrix}$ (62)
Action Execution $P\xrightarrow{Controllers}\left[\delta_{f},a\right]\rightarrow ECU$
V-B3 RAG-Enhanced Memory Module
Given that our 8B-parameter VLM lacks the extensive reasoning capabilities of larger, 100B-200B parameter models, we employ a RAG-based approach [lewis2020retrieval] and integrate it to enhance reasoning and enable human feedback learning. This system is built upon the Chroma vector database [chroma_ai-native_2023], enabling efficient storage history interactions and retrieval of similar driving scenarios.
The memory module is uniquely created for each user, ensuring a personalized driving experience follows individual preferences and patterns. It stores driving-related information in a structured format comprising commands paired with corresponding context tuples:
$\{I-(I,D,P,F)\}$ (63)
When processing a new driving scenario, the instruction $\boldsymbol{I}$ is used for similarity matching to retrieve the top-k similar prior situations. The associated data values are then sent to the VLM, enhancing decision-making with relevant context and supporting personalization. This RAG-enhanced memory enables the VLM to handle similar situations with greater accuracy and consistency, improving the vehicle’s responsiveness to unique user preferences and enhancing the overall driving experience.
V-B4 Multi-Controller Joint Motion Control
As shown in Fig. 9, we implement a decoupled control strategy that separates lateral and longitudinal vehicle motion control. The lateral control is handled by MPC calculating the front steering angle $\delta_{f}$, while longitudinal control is managed through a PID controller calculating longitudinal acceleration $\alpha$. The motion planning module in the upper layer provides trajectories consisting of waypoints, which our motion control system tracks. The calculated $\alpha$ and $\delta_{f}$ are then transmitted to the drive-by-wire system developed by AutonomousStuff [autostuff] for precise control of throttle, braking, and steering.
For the longitudinal control, the PID controller calculates the $\alpha$ for each time step $\Delta t$ to minimize the velocity error $e_{v}$, which is the difference between the current velocity $V_{\text{current}}$ and the desired velocity $V_{\text{ref}}$.
$\alpha(t)=K_{p}e_{v}(t)+K_{i}\sum_{i=0}^{t}e_{v}(i)\Delta t+K_{d}\frac{\Delta e_{v}(t)}{\Delta t}$ (64)
where $K_{p}$, $K_{i}$, and $K_{d}$ are the proportional terms, integration terms, and derivative terms that will be contained in the action matrix generated by our VLM.
For the lateral control, our MPC approach utilizes a linear vehicle dynamics model [snider2009automatic] to predict future states and optimize the front steering angle, $\delta_{f}$, over a finite prediction horizon. With the prediction model[snider2009automatic], the control increment is then obtained by solving a Quadratic Program (QP) [bemporad2000explicit] to minimize the cost function $J$ in the MPC:
$J=E^{T}QE+\Delta_{f}^{T}R\Delta_{f}$ (65)
where $E$ is the predicted state calculated by the prediction model, and $\Delta_{f}$ is the future control input. The $Q$ and $R$ are weighting matrices that penalize tracking state deviations and control effort. Our VLM generates the action matrix that primarily considers three key components: the weight for lateral error ($W_{l}\in Q$), the weight for heading error ($W_{h}\in Q$), and the weight for the squared terms of speed and steering inputs ($W_{s}\in R$). These weighting factors are selected as they demonstrate the most significant impact on lateral control performance.
V-B5 Efficient On-Board VLM Module
We generate a dataset of 10,000 image-instruction pairs, each labeled with the desired action, to create a comprehensive training set for fine-tuning our on-board VLM. This VLM is based on the Qwen-VL architecture [bai2023qwenvlversatilevisionlanguagemodel], which we fine-tune using the Low-Rank Adaptation (LoRA) method [hu2021loralowrankadaptationlarge] (a type of Parameter Efficient Fine-Tuning (PEFT) [xu2023parameterefficientfinetuningmethodspretrained]), enabling significant customization while preserving computational efficiency. To optimize for on-board deployment, we apply 4-bit Activation-Aware Weight Quantization (AWQ) [lin2024awqactivationawareweightquantization], compressing the VLM to increase inference speed without sacrificing too much accuracy. This combination of techniques ensures a responsive, on-board VLM suited to real-time response.
Dataset Collection
We develop a specialized training dataset to fine-tune the Qwen-VL model [bai2023qwenvlversatilevisionlanguagemodel], consisting of 10,000 semi-human-annotated image-text pairs. Each image, representing a traffic scene sourced from the NuScenes dataset [caesar_nuscenes_2020], which includes numerous diverse traffic scenarios, is paired with a human-provided instruction and a corresponding action label in the form of a controller action matrix, guiding the model’s response in different traffic scenarios.
The human instructions are also very diverse, ranging from explicit commands like “speed up” to more implicit cues such as “I am in an urgent situation.” This diversity allows the VLM to interpret both clear and vague inputs, improving its ability to understand complex human intentions. To enhance the model’s responsiveness to different driving styles, we annotate each image with three different instruction types—aggressive, moderate, and conservative—each paired with a corresponding action. This multi-faceted approach ensures that the VLM can adapt its behavior to match various driving styles, enabling it to respond flexibly and contextually across diverse traffic conditions.
LoRA Finetune
We apply the LoRA method to fine-tune our Qwen-VL model. LoRA works by freezing the pre-trained model weights and introducing trainable, low-rank decomposition matrices into each layer of the Transformer architecture. This approach significantly reduces the number of trainable parameters required, making fine-tuning more efficient.
The fine-tuning process for our VLM is conducted on a cluster of four NVIDIA A100 GPUs, each equipped with 40GB of memory. The model is trained over five epochs with a per-device batch size of two for training and one for validation, using a learning rate of 1e-5. Additionally, we implemented gradient accumulation with eight steps, allowing for effective larger batch processing. This setup enables the entire training process to be completed in approximately five hours, ensuring both accuracy and efficiency in model performance.
Compression and On-Board Deployment of VLM
AWQ [lin2024awqactivationawareweightquantization] is a hardware-friendly technique for low-bit, weight-only quantization, specifically designed for VLM. AWQ minimizes quantization error by identifying the 1% salient weights, which are then scaled using an equivalent transformation to preserve their precision. We apply AWQ to quantize our model to INT4, achieving improved quantization performance suited for on-board deployment. Additionally, we utilize the LMDeploy toolkit [2023lmdeploy] to optimize inference time. This enhancement is made possible through features such as persistent batching, blocked KV cache, tensor parallelism, and optimized CUDA kernels, all of which contribute to high-performance, low-latency operation.
V-C Real-World Experiment
To comprehensively evaluate our system’s performance, we conduct a series of experiments assessing its ability to provide safe, comfortable, reliable, and personalized driving experiences. We employ multiple evaluation metrics: a driving score to measure driving performance, including safety, comfort, and alignment with environmental conditions and human instructions; takeover frequency to assess personalization capabilities. Additionally, we perform an ablation study to examine the effectiveness of the memory module. The demo for our real-vehicle experiment is publicly available [^3] Real-vehicle experiment demo video with on-board VLM: https://youtu.be/lxzXhTasSdMover.
TABLE IX: Driving Performance Validation. $\downarrow$: Lower Values are Better. $\uparrow$: Higher Values are Better.
[…]
Time to
Collision($s$$\uparrow$
$SV_{x}$
($m^{2}/s^{2}$$\downarrow$
$SV_{y}$
($10^{-2}m^{2}/s^{2}$$\downarrow$
$|\bar{\alpha_{x}}|$
($m/s^{2}$$\downarrow$
$|\bar{J_{x}}|$
($m/s^{3}$$\downarrow$
$|\bar{\alpha_{y}}|$
($10^{-1}m/s^{2}$$\downarrow$
$|\bar{J_{y}}|$
($m/s^{3}$$\downarrow$
VLM
Latency($s$$\downarrow$
Command
[…]
Scenario
[…]
Acceleration Baseline 2.44 28.8 0.36 0.78 3.27 0.46 0.44 - 92.0 60.0 75.6
GPT-4o 2.52 30.0 0.39 0.83 3.40 0.52 0.52 5.82 92.9 71.3 76.4
Ours 2.46 30.8 0.39 0.81 3.07 0.53 0.81 1.98 96.3 60.9 76.5
Lane Change Baseline 2.44 3.91 1.65 0.37 3.14 0.88 1.01 - 88.5 60.0 74.5
GPT-4o 2.71 3.88 2.23 0.53 4.38 1.13 1.39 4.84 90.4 88.6 78.4
Ours 2.15 4.07 2.15 0.41 3.35 0.98 1.02 1.83 92.2 71.9 77.5
Left Turn Baseline - 1.12 7.52 0.22 1.81 1.29 1.36 - 88.0 60.0 70.4
GPT-4o - 0.93 11.5 0.29 2.75 2.11 2.40 5.23 91.3 85.0 71.4
Ours - 0.94 6.74 0.19 1.67 1.33 1.32 1.64 90.2 67.8 74.4
Figure 10: Overview of the experiment track.
V-C1 Experiment Setup
Test Track and Participants
The field experiments [^4] The experiments conducted in this study satisfy all local traffic guidelines and guarantee the safety of the participants. A human always sits in the driver’s seat of the autonomous vehicle to monitor its status and get ready to take over. aim at validating the real-world performance of our personalized motion control system. We include three types of driving behaviors—accelerating, lane changing, and turning—to comprehensively validate control over steering, throttle, and braking. An overview of the test track and driving behaviors is shown in Fig. 10. For both acceleration and lane change scenarios, a lead vehicle is positioned 30 $m$ ahead of the ego vehicle, accelerating from static to 45 $km/h$ with an acceleration of 1.26 $m/s^{2}$. In the acceleration scenario, the ego vehicle accelerates from a complete stop to reach 50 $km/h$. In the lane change scenario, the ego vehicle maintains 50 $km/h$ while overtaking the lead vehicle. For the intersection turning scenario, the ego vehicle navigates a curve with a radius of 23.89 $m$ at a constant speed of 30 $km/h$.
V-C2 System Driving Performance
To showcase our system’s driving performance, we conduct comparative experiments against two systems: a baseline system using pre-defined controller parameters for general safety and comfort and a system utilizing GPT-4o with few-shot learning. We evaluate across three primary meta-driving scenarios (acceleration, lane changing, and turning), with each scenario tested under ten different commands and five weather conditions (sunny, rain, fog, snow, and night) to test the model’s vision understanding as shown in Fig. 11.
Evaluate Metrics
The models are then assessed based on four key aspects—safety, comfort, time efficiency, and alignment, as shown in Tab. IX. An overall driving score $S$ is then calculated as a weighted sum of the individual metric scores, denoted as:
$S=\sum{w_{k}\cdot S_{k}}$ (66)
where $k$ includes all ten metrics: Time to Collision ($\tau$), longitudinal and lateral speed variance ($SV_{x}$ and $SV_{y}$), lateral and longitudinal mean absolute acceleration ($|\bar{\alpha_{x}}|$ and $|\bar{\alpha_{y}}|$), lateral and longitudinal mean absolute jerk ($|\bar{J_{x}}|$ and $|\bar{J_{y}}|$), VLM Latency, Command Alignment and Scenario Alignment. All the scores $S_{k}$ range from 0 to 100, while the weights of scores $w_{k}$ are empirically tuned for each driving scenario. For instance, longitudinal parameters have higher weights in acceleration scenarios, while lateral parameters are weighted more heavily in turning scenarios. The safety and comfort metrics are followed as defined in Sec. IV-C3.
Figure 11: Sample vision inputs from different weather scenarios.
The alignment evaluation consists of two aspects. Command Alignment quantifies the deviation between the model output and the expected parameter range, calculated as a weighted average across six control parameters. For each parameter, we establish three ranges based on past experiments, corresponding to aggressive, conservative, or moderate driving styles. Taking the PID controller’s proportional parameter $K_{p}$ as an example, the score is calculated as:
$S_{K_{p}}^{*}=\begin{cases}\frac{100(K_{p}-K_{p,\min})}{K_{p,\text{lower}}-K_{p,\min}},&K_{p}\in[K_{p,\min},K_{p,\text{lower}}),\\[10.0pt] 100,&K_{p}\in[K_{p,\text{lower}},K_{p,\text{upper}}),\\[10.0pt] \frac{100(K_{p,\max}-K_{p})}{K_{p,\max}-K_{p,\text{upper}}},&K_{p}\in[K_{p,\text{upper}},K_{p,\max}],\\[10.0pt] 0,&\text{else}.\end{cases}$
where $K_{p,\min}$ and $K_{p,\max}$ are the minimum and maximum overall parameter range obtained through experiments, while $K_{p,\text{lower}}$ and $K_{p,\text{upper}}$ are determined by the command intention labeled by human experts. The scenario alignment score computes whether the model can capture details of the scene through vision input and act more aggressively or conservatively based on the current condition. It is calculated by tallying the percentage of instances where the model gives more conservative parameter sets in adverse weather conditions compared to the sunny clear scenarios.
Result
As shown in Tab. IX, the command alignment score of our VLM model is similar to or greater than GPT-4o, showing our model has high reasoning capabilities regarding the personalization of control parameters. As for the scenario alignment, our model shows significantly better performance than baseline conditions in lane changing and left turn scenarios but a score very similar to the baseline scenario. We think this is mostly due to the model considering acceleration on a clear straight road in less dangerous situations and thus does not act too conservatively. In terms of the overall Driving Score, our model also surpasses the performance of baseline models and even GPT-4o in some scenarios, indicating our model provides the most well-rounded driving experience.
V-C3 Human-in-the-Loop Validation
Figure 12: Takeover rate comparison between the baseline and our method.
This subsection evaluates the effectiveness of our method in reducing takeover rates compared to the baseline system. The baseline consists of the traditional autonomous vehicle controller with default settings, where two unified controllers manage vehicle operations on longitudinal and lateral respectively. We compare this conventional approach against our VLM-based adaptive motion control system. Throughout the experiments, participants could provide explicit instructions or implicit preferences/feedback with varying degrees of directness, prompting the system to make corresponding adjustments. The instructions were categorized into three levels of directness, as defined in Tab. V.
Every participant is supposed to provide at least five instructions for each scenario. For each instruction-scenario pair, participants completed two trips - one with the baseline system and one with our VLM solution. To ensure unbiased evaluation, participants are not informed which system they are using during each trip. We use the takeover rate as our primary metric to evaluate the system’s ability to provide personalized driving experiences.
The results demonstrate that our VLM-based method consistently outperforms the baseline system across all three levels of instruction directness. With Level 1 (direct) instructions, our method achieves a 5.56% takeover rate versus the baseline’s 19.44%, representing a 71.4% reduction. For Level 2 (moderately direct) instructions, the rates are 6.06% versus 12.12%, showing a 50% improvement. Most notably, for Level 3 (indirect) instructions, our method achieves 8.33% versus the baseline’s 36.11%, marking a 76.9% reduction in takeover rate. The result demonstrates our system’s better capability in interpreting and acting based on user preferences, regardless of how explicitly they are communicated.
V-C4 Ablation Study on Memory Module
Figure 13: Effectiveness of memory modules in takeover rates
To further validate the effectiveness of our RAG-based Memory Module (MM), we conduct an ablation study with three human participants, comparing three configurations: our complete system with MM, the system without MM (W/O MM), and the baseline. The results demonstrate the significant impact of the MM on reducing takeover rates.
As shown in Fig. 13, With three participants, our complete system with the memory module achieves the lowest average takeover rate of 6.67%. When we remove the memory module while keeping other components the same, the average takeover rate increases substantially to 24.44%. The baseline system performs worst with a 44.44% average takeover rate. These results indicate a 72.7% reduction in takeover rate when adding the memory module to our base architecture and an 85% overall reduction compared to the baseline. This significant improvement suggests that the RAG-based memory module plays a crucial role in maintaining personalized vehicle control by effectively utilizing historical interactions and user preferences.
VI Challenges
VI-A Latency
Latency presents a fundamental challenge in deploying LLMs for autonomous driving applications. While traditional autonomous driving systems operate with millisecond-level response times, the integration of LLMs introduces significant latency concerns that must be carefully addressed. Our experimental work with cloud-based LLM APIs revealed multiple layers of latency challenges that impact system reliability.
Network latency forms the first critical bottleneck. Our real-world experiments under stable 4G-LTE conditions showed a baseline communication latency of approximately 300ms - already an order of magnitude higher than typical autonomous driving control loops. This network dependency introduces additional vulnerabilities, as real-world driving conditions often involve varying network quality, dead zones, or connection interruptions that could compromise system reliability.
The LLM processing itself introduces substantial additional latency. Our empirical results with text-based GPT-4 [openai_gpt-4_2023] or onboard 9B VLM showed end-to-end latencies ranging from 1.2 to 1.8 seconds, while the visual-capable GPT-4o [openai2024gpt4ocard] exhibited even higher latencies around 5 seconds (assuming 300ms network latency). These latencies pose significant challenges for real-time decision-making in autonomous driving scenarios. While such response times might be acceptable for high-level decision-making or handling non-critical user requests, they become problematic for any safety-critical functions.
The latency challenge is further complicated by the variable nature of LLM response times. Input complexity, prompt length, and required reasoning depth can all significantly impact processing time. This variability makes it difficult to guarantee consistent response times - a crucial requirement for safety-critical systems. Additionally, the system must manage the trade-off between prompt effectiveness and latency, as more detailed prompts that might lead to better responses also increase processing time.
These latency constraints limit the role LLMs can play in autonomous driving architectures. Critical path functions requiring real-time responses (collision avoidance, emergency braking, immediate obstacle response) may still remain handled by traditional, low-latency control systems. This creates a complex architectural challenge: how to effectively partition decision-making responsibilities between high-latency LLM components and low-latency control systems while maintaining system safety and reliability.
VI-B Deployment
Another important consideration is the deployment of onboard LLMs in vehicles. While cloud-based LLM APIs offer flexibility and scalability, they rely on stable network connections, which may not always be available in all driving scenarios. To mitigate this issue, researchers are exploring the possibility of integrating LLMs directly into the vehicle’s onboard systems. However, the deployment of LLMs in autonomous vehicles faces significant computational resource limitations. Automotive hardware typically has lower computational capabilities compared to development systems, creating a substantial gap between model requirements and available resources [cui2024onboardvisionlanguagemodelspersonalized]. This constraint affects all aspects of system performance, from basic perception tasks to complex decision-making processes. The challenge is particularly obvious for LLMs, which require substantial GPU computing power that often exceeds typical automotive hardware capabilities.
Additionally, researchers can focus on optimizing LLMs specifically for deployment in vehicles. This involves streamlining the models to reduce computational requirements, enabling them to run efficiently on the limited hardware resources available in automotive systems. Model distillation, which transfers knowledge from large teacher networks, such as multimodal LLMs, to smaller student architectures, is a promising yet underexplored direction in the autonomous driving domain for addressing computational constraints [sreenivas2024llm]. Similarly, plug-and-play approaches offer another compelling avenue for integrating foundation model capabilities into on-device autonomous driving networks without adding additional computational complexity [ma2023lampilot].
Researchers must find a balance between model complexity and hardware constraints to ensure practical deployment in vehicles. As developments in hardware and model optimization continue, the feasibility of integrating powerful onboard LLMs directly into vehicles will likely increase, opening up new possibilities for enhancing the capabilities of autonomous driving systems.
VI-C Security and Privacy
The integration of LLMs into autonomous driving systems introduces significant security and privacy vulnerabilities that require careful consideration. While traditional autonomous vehicles already face cybersecurity challenges, the addition of LLMs creates new attack surfaces due to their natural language processing capabilities and complex decision-making processes.
A fundamental security concern lies in adversarial attacks targeting the LLM’s language understanding capabilities [goodfellow2014explaining, carlini2017towards]. Unlike traditional autonomous systems where inputs are structured and validated, LLMs must process natural language commands that could contain malicious intent. For instance, carefully crafted commands could potentially trigger the model’s semantic understanding to trigger unsafe driving behaviors or override safety protocols. These attacks are particularly challenging to defend against because they operate at the semantic level, where distinguishing between approprate and malicious commands becomes highly complex.
The risk of data poisoning and backdoor attacks [biggio2012poisoning, liu2018trojaning] presents another critical challenge, particularly given the continuous learning and adaptation capabilities of LLM-based systems. Attackers could potentially inject malicious training examples that create hidden triggers in the model’s language understanding. For example, specific phrases or conversation patterns could be embedded as triggers that cause the system to deviate from safe operation. The detection of such backdoors becomes increasingly difficult as the model’s decision-making process grows more sophisticated through personalization and learning from user interactions.
Prompt injection attacks [bagdasaryan2023ab] represent a unique threat in the context of autonomous driving. Since LLMs process natural language inputs to generate driving decisions and responses, attackers could potentially embed harmful instructions within seemingly innocent conversations. For instance, a casual conversation about route preferences could contain hidden commands that manipulate the vehicle’s behavior in subtle but dangerous ways.
Privacy concerns extend beyond traditional vehicular data protection. LLM-based autonomous systems necessarily process and retain extensive personal data, including individual communication patterns and preferences, regular routes and destinations, context-specific behaviors and habits, voice characteristics and linguistic patterns, emergency response preferences, and medical information. This rich personal data, combined with the system’s cloud-based architecture, creates significant privacy risks. Traditional privacy-preserving techniques like differential privacy [abadi2016deep] face new challenges in this context, as they must balance data protection with the need for personalized and context-aware responses.
VI-D Safety
When LLMs are integrated into autonomous driving systems, safety is always the first priority. In our proposed LLM4AD concept, the LLMs themselves do not directly handle safety requirements because of the potential for hallucinations in LLM outputs and the reaction time required. Hallucinations could be dangerous in safety-critical situations, and LLMs may not react quickly enough to sudden events. In our solutions, we employ a rule-based safety guarantee mechanism where LLM outputs must fall within predefined safe ranges - if outputs exceed these bounds or do not match the output format, execution is automatically blocked.
However, the challenge is that ensuring system safety and reliability requires extensive validation across a wide range of operational conditions. This includes testing system responses to edge cases, verifying behavior in degraded operating conditions [he2010single, zhu2024mwformer, 10310160, li2024light], and validating the effectiveness of fallback mechanisms. The challenge extends to maintaining clear audit trails of system decisions and ensuring alignment with different regulatory requirements. This is particularly critical for systems implementing LLM-based systems, where behavior must consistently align with defined ethical guidelines while maintaining safety guarantees.
VI-E Trust and Transparency
A critical challenge in integrating LLMs into autonomous driving systems lies in establishing genuine trust and transparency between human users and the autonomous vehicle. While LLMs can generate natural language explanations for vehicle decisions [marcu2024lingoqavisualquestionanswering], several fundamental challenges emerge:
First, there is the risk of explanation mismatch - the LLM might generate plausible-sounding explanations that do not actually match the true decision-making process of the underlying autonomous system. This could create a false sense of transparency and potentially dangerous misconceptions about the vehicle’s capabilities and limitations. When the LLM generates both executable policies and natural language reasoning, we must ensure perfect alignment between the two to avoid misleading explanations.
Apart from that, the real-time nature of autonomous driving poses significant challenges for LLM-based explanations. The system must balance between providing detailed, comprehensible explanations and maintaining rapid response times for critical driving decisions. Too much detail could overwhelm users in time-sensitive situations, while oversimplified explanations might omit crucial context.
Additionally, the black-box nature of LLMs themselves raises questions about the trustworthiness of their explanations. Users might question whether they can trust explanations generated by a system that operates in ways that are not fully transparent. This creates a meta-level transparency challenge that needs to be carefully addressed in the system design.
VI-F Personalization
Although one of the primary focuses of our research is integrating LLMs to achieve a higher level of personalization in autonomous vehicles by learning and adapting to individual drivers’ habits, preferences, and communication styles over time, it also introduces significant technical and safety challenges.
A critical challenge lies in the interpretation of natural language commands and emotional states. While LLMs can process natural language, accurately inferring the driver’s true intent and appropriate response is non-trivial. For instance, if a driver expresses urgency with “I am on my way to the hospital,” the system must carefully balance between accommodating the emergency situation and maintaining safe driving practices. Simply adopting a more aggressive driving style could introduce unnecessary risks, particularly in complex traffic conditions where the safest course of action might conflict with the inferred urgency.
The long-term learning and adaptation process also raises important questions about system boundaries and safety constraints. As the system builds a personalized profile over time, it must maintain clear limits on which aspects of driving behavior can be personalized and which must remain constrained by safety parameters. This creates a complex balance between personalization and standardized safety requirements, particularly in scenarios where learned preferences might conflict with optimal safety decisions.
VII Conclusions
In this paper, we presented a comprehensive exploration of LLM4AD, the application of Large Language Models (LLMs) for autonomous driving. Our research highlighted the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making. By leveraging the natural language understanding and reasoning abilities of LLMs, we demonstrated their capacity to enhance the safety, efficiency, and user experience of autonomous vehicles. Moreover, we introduced a comprehensive benchmark specifically designed to evaluate the instruction-following abilities of LLMs within the autonomous driving domain. Our experiments, conducted on real-world vehicle platforms, showed promising results, validating the effectiveness of our proposed LLM-based approaches. These findings underscored the significant impact that LLMs can have on the future of autonomous driving technology, making it safer, smarter, and more accessible for all.
[…]
[Uncaptioned image] Can Cui (Student Member, IEEE) received his M.S. degree in Electrical and Computer Engineering from the University of Michigan, Ann Arbor in 2022. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University. His research pursuits encompass automated driving, Advanced Driver-Assistance Systems (ADAS), applied machine learning, controls, and digital twin.
[Uncaptioned image] Yunsheng Ma (Student Member, IEEE) received his M.S. degree in Computer Science from New York University in 2022. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on autonomous driving, embodied AI, and digital twins. He has publications in top-tier conferences and journals, including CVPR, AAAI, ITSC, and IEEE Transactions on Intelligent Vehicles.
[Uncaptioned image] Zichong Yang received his M.S. degree in Mechanical Engineering from Carnegie Mellon University in 2023. He is now pursuing a Ph.D. degree in the College of Engineering at Purdue University, with a research focus on computer vision, human pose estimation, and human trajectory prediction.
[Uncaptioned image] Yupeng Zhou received his B.S. degree in Mechanical Engineering from South China University of Technology in 2022. He is now an M.S. student in the College of Engineering at Purdue University, with a research focus on applied robotics, autonomous vehicles, and controls.
[Uncaptioned image] Peiran Liu received a B.S. degree in Information Systems from the University of Florida in 2023. He is currently pursuing his Ph.D. degree in the College of Engineering at Purdue University. His research interests include computer vision, human behavior predictions, and digital twins.
[Uncaptioned image] Juanwu Lu (Student Member, IEEE) received an M.S. degree in Civil Engineering from the University of California, Berkeley in 2022. He is currently pursuing his Ph.D. degree in the College of Engineering at Purdue University. His research interests include statistical machine learning and its application for generalizable and interpretable microscopic traffic modeling.
[Uncaptioned image] Lingxi Li (Senior Member, IEEE) received his Ph.D. degree in Electrical and Computer Engineering from the University of Illinois at Urbana–Champaign in 2008. He is a professor in Electrical and Computer Engineering at Purdue University. His research interests include complex systems, intelligent transportation systems, connected and automated vehicles, discrete event dynamic systems, human-machine interaction, digital twins, and parallel intelligence.
[Uncaptioned image] Yaobin Chen (Senior Member, IEEE) received his Ph.D. degree in Electrical and Computer Engineering from Rensselaer Polytechnic Institute in 1988. He is a professor in Electrical and Computer Engineering at Purdue University. His areas of expertise include automated driving and autonomous/connected vehicles, vehicle active safety systems, EV and HEV, intelligent controls and robotics, and advanced energy and power systems.
[Uncaptioned image] Jitesh Panchal received his Ph.D. degree in Mechanical Engineering from Georgia Institute of Technology in 2005. He is a professor in Mechanical Engineering at Purdue University. His areas of expertise include computational design of socio-technical systems, secure design and manufacturing, engineering design by self-organized virtual communities, and integrated products and materials design.
[Uncaptioned image] Amr Abdelraouf (Member, IEEE) received his Ph.D. degree in Transportation Engineering from the University of Central Florida in 2022. He is a research scientist at Toyota Motor North America, InfoTech Labs. His research interests include applied AI/ML, advanced driver assistance systems, connected and automated vehicles, intelligent transportation systems, and digital twins.
[Uncaptioned image] Rohit Gupta received the Ph.D. degree in Mechanical Engineering from the University of California at Santa Barbara. He is currently a principal researcher at Toyota Motor North America, InfoTech Labs. Prior to joining Toyota, he worked in various industries: semiconductors, telecommunications, finance, and software. His research interests include cloud and edge computing architectures, connected and autonomous vehicles, AI/ML, “Digital Twin” of human-vehicle interactions, and high-speed mobility.
[Uncaptioned image] Kyungtae (KT) Han (Senior Member, IEEE) received the Ph.D. degree in Electrical and Computer Engineering from The University of Texas at Austin in 2006. He is currently a senior principal scientist at Toyota Motor North America, InfoTech Labs. Prior to joining Toyota, Dr. Han was a Research Scientist at Intel Labs and a Director at Locix Inc. His research interests include cyber-physical systems, connected and automated vehicle techniques, and intelligent transportation systems.
[Uncaptioned image] Ziran Wang (Senior Member, IEEE) received his Ph.D. degree in Mechanical Engineering from the University of California, Riverside in 2019. He is an assistant professor in the College of Engineering at Purdue University, and was a principal researcher at Toyota Motor North America R&D. His research focuses on automated driving, human-autonomy teaming, and digital twins. He is the Founding Chair of the IEEE Technical Committee on Internet of Things in Intelligent Transportation Systems.